### v1

# ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT

Abstract.
Technical question and answering (Q&A) sites such as Stack Overflow have become an important source for software developers to seek knowledge. However, code snippets on Q&A sites are usually uncompilable and semantically incomplete for compilation due to unresolved types and missing dependent libraries, which raises the obstacle for users to reuse or analyze Q&A code snippets. Prior approaches either are not designed for synthesizing compilable code or suffer from a low compilation success rate. To address this problem, we propose ZS4C, a lightweight approach to perform zero-shot synthesis of compilable code from incomplete code snippets using Large Language Model (LLM). ZS4C operates in two stages. In the first stage, ZS4C utilizes an LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template. In the second stage, ZS4C fixes compilation errors caused by incorrect import statements and syntax errors through collaborative work between ChatGPT and a compiler. We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO against the SOTA approach SnR. Compared with SnR, ZS4C improves the compilation rate from 63% to 87.6%, with a 39.3% improvement. On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1.
imcomplete code snippets, Large Language Model, Program synthesis, API inference, Prompt Engineering, ChatGPT
1. Introduction
Technical question and answering sites (Q&A) such as Stack Overflow have become an important source for software developers to seek knowledge, including code snippets that demonstrate API usage. A recent study revealed that 15 million posts on Stack Overflow have source code snippets, forming a vast codebase for developers to reuse (Wu et al., 2019). Recent surveys reported that developers search for code snippets in Q&A sites frequently and extensively (Sadowski et al., 2015; Wu et al., 2019; Xia et al., 2017).
However, the majority of the code snippets on Q&A sites are uncompilable, which raises the obstacle for users to reuse or analyze these code snippets (Terragni et al., 2016; Rubei et al., 2020). Those code snippets are usually uncompilable and semantically incomplete for compilation due to unresolved types (e.g., what is the fully qualified name (FQN) for a type DateFormat) and missing dependent libraries (e.g., which library to import for DataFormat) (Terragni et al., 2016; Mishne et al., 2012; Dong et al., 2022). Although missing such implementation details could be synthesized manually, it is tedious and often requires substantial familiarity with various libraries from developers. Furthermore, such manual synthesis effort is, however, not scalable for software engineering research based on Q&A sites’ code snippets (Zhang et al., 2021, 2019; Gao et al., 2015; Mishne et al., 2012). A technique to automatically repair code snippets by resolving the missing types and dependent libraries for compilation will significantly save developers time and facilitate the research based on such code snippets.
To address this issue, various approaches have been proposed (Terragni et al., 2016; Dong et al., 2022; Saifullah et al., 2019; Subramanian et al., 2014; Phan et al., 2018). For instance, CSNIPPEx synthesizes compilable code by inferring dependency libraries for code snippets based on a set of hand-crafted heuristics (Terragni et al., 2016). Baker (Subramanian et al., 2014) and SnR (Dong et al., 2022) extract constraints from the abstract syntax tree (AST) of code snippets and use constraint solving to infer the FQN for missing types and their corresponding dependent libraries. StatType (Phan et al., 2018) and Coster (Saifullah et al., 2019) build statistical models from existing compilable source code to predict FQN for reresolved types. However, Coster and StatTyper are not designed directly for synthesizing compilable code from code snippets, instead, it designed for FQN inference. For the approaches (e.g., Baker, CSNIPPEx, and SnR) that are designed for synthesizing compilable code, their reported compilation success rates are lingering below 80%, it’s evident that there’s room for developing new tools to improve compilation rates. Thus, we aim to explore the possibility of designing a lightweight tool, capable of addressing compilation errors in incomplete code snippets, without relying on pre-defined rules and program analysis.
Large language models, such as ChatGPT, have demonstrated their effectiveness in various software engineering tasks (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023), serving as a knowledge base for consultation. The two key steps for synthesizing compilable code from incomplete code snippets, inferring dependent libraries, and fixing potential compilation errors naturally align with the strength of LLM.
In this study, we introduce ZS4C, an approach designed to perform zero-shot synthesis of compilable code from incomplete code snippets using LLM. ZS4C operates in two stages. In the first stage, ZS4C utilizes LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template. We apply the widely adopted self-consistency method to improve the robustness of the inference. In the second stage, ZS4C employs the Multi-Agent Conversation framework (Wu et al., 2023) to fix compilation errors caused by incorrect import statements and syntax errors. Specifically, ZS4C uses a compiler as an agent to provide feedback, i.e., compilation error logs, which serve as additional context for ChatGPT to correct the errors. Compared with prior approaches, ZS4C is lightweight and does not require any training data, any pre-defined rules or heuristics, or program analysis like constructing AST.
We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO consisting of 267 code snippets manually collected from Stack Overflow posts (Phan et al., 2018) against the SOTA approach SnR. ZS4C improves the compilation rate of SnR, from 63% to 87.6%, with a 39.3% improvement. On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1 score. Our ablation analysis shows that both components in ZS4C play a crucial role in the effectiveness of ZS4C. For instance, by employing the component conversation between a compiler and ChatGPT, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement. Our parameter analysis reveals that ZS4C is not sensitive to parameters. Lastly, we conducted a failure analysis and our results sheds light on the limitations of using LLMs for synthesizing code, such as hallucination (e.g., fake FQN, and partially correct FQN) and unexpected code modification from LLMs.
This paper makes the following contributions:
  • We propose ZS4C, a novel zero-shot LLM-based approach to automatically synthesize compilable code for code snippets, which employs the multi-agent conversation framework to boost the successful rate of synthesizing compilable code for code snippets.
  • We conducted extensive experiments to demonstrate the effectiveness of ZS4C compared with SOTA approach, and other baselines that mimic how developers use ChatGPT for resolving compilation issues for incomplete code snippets.
  • We shed light on the limitation of using LLMs for the code synthesis task.
  • We made our code and dataset publicly available (dat, 2023) to encourage future research on studying partial code synthesis and related approaches.
2. Background and related work
2.1. Automated Synthesis of Compilable Code from Stack Overflow
The task of automatically synthesizing compilable code snippets for incomplete code snippets on Q&A is first introduced by Terragni et al. (Terragni et al., 2016). They developed an Eclipse plug-in named CSNIPPEx for this task. The tool consists of three main components: C-Units Inference, Dependencies Resolver, and Code Completer. The C-Units Inference component is for resolving missing type declarations. The Dependencies Resolver component automatically identifies the necessary external libraries in their FQNs (fully qualified names, e.g., org.library1.sub.C) [^1] Without fully qualified names, naming ambiguities can easily arise, as many simple class names across different libraries collide with one another. based on a set of heuristics and generates import declarations. Finally, the Code Completer addresses issues like syntax errors, typos, and missing variable declarations by using Eclipse Quick Fix. A follow-up tool called APIzation was developed on top of CSNIPPEx to wrap up a code snippet and generate a reusable API with proper parameters, return statements, and missing external libraries (Terragni and Salza, 2021). ZS4C draws inspiration from CSNIPPEx and consists of two components: one dedicated to identifying necessary library dependencies, and the other focused on iteratively improving the code by compiling and correcting errors. One main difference is that we utilize carefully designed prompts to leverage the power of LLM (ChatGPT) for the two sub-tasks and no pre-defined heuristics and program analysis (i.e., AST construction) are required.
Another work from Dong et al (Dong et al., 2022), in which they proposed a tool named SnR, that can automatically infer FQNs, recommend libraries, and create import statements for code snippets, eventually making code compilable. SnR is a constraint-based approach. It first analyzes the incomplete code snippet and extracts constraints that capture the relationships between different elements in the code. Next, SnR utilizes a knowledge base built from a collection of real Java libraries which contains information about classes, methods, and dependencies. The extracted constraints are then solved using a Datalog solver called Soufflé (Dong et al., 2022). This solver uses the knowledge base to find solutions that satisfy the constraints. Once the constraints are solved, SnR transforms the solutions into import statements. These import statements are added to the code snippet to resolve missing dependencies. They have compared their tool with the state-of-the-art FQN inference approach named COSTER (Saifullah et al., 2019). Experiments on a benchmark of 267 code snippets from Stack Overflow show that SnR correctly infers 91.0% of the import statements, which makes 73.8% of the snippets compile, compared to 36.0% of the import statements and 9.0% of the snippets by COSTER (Saifullah et al., 2019). Given these results, we consider SnR as our baseline approach.
One of the key tasks in the automated synthesis of compilable code snippets is to identify the fully qualified name (FQN) for APIs (FQN inference(Huang et al., 2022). Specifically, identifying the types of API elements in incomplete code can help identify the dependent libraries. Existing studies addressed FQN inference in different ways: Baker extracts constraints from code snippets and uses a naive constraint solving algorithm to infer FQNs (Subramanian et al., 2014); StatType formulates the task as a phrase-based machine translation task - translating the code snippet into a target sentence that contains the FQNs of the API elements (Phan et al., 2018) and builds a statistical model from existing compilable source code; COSTER (Saifullah et al., 2019) is another statistical model-based approach, it calculates the likelihood of appearing context tokens and the FQN of each API element, and then ranks the FQNs based on three different scores: likelihood, context similarity, and name similarity scores; Huang et al. (Huang et al., 2022) treats FQN inference as clozy-style fill-in-blank language task. They utilized the pre-trained CodeBERT model (Feng et al., 2020) and fine-tuned it using prompt learning with FQN-annotated code examples, sampled from target libraries. The prompt learning stimulates the MLM to recognize FQN syntax and usage patterns. In our approach, we do not directly perform FQN inference, instead, we skip it and directly focus on import statement inference.
2.2. Using LLMs for Software Engineering Tasks
With the breakthrough of LLMs in the field of natural language processing, there’s a burgeoning interest in investigating the potential of LLMs to enhance software engineering tasks. Researchers have pre-trained LLMs on publicly available code, resulting in the development of specialized code LLMs like CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), Codex (Chen et al., 2021), and CodeGen (Nijkamp et al., 2022). In addition to these pre-trained code models, another research trajectory focuses on optimizing the application of LLMs to software engineering tasks such as code generation (Liu et al., [n. d.]), program repair (Wei et al., 2023), testing (Schäfer et al., 2023), and beyond. Given the remarkable capability of LLMs to learn from zero or minimal examples, there’s a growing exploration into how in-context learning and prompt engineering can elevate LLM-based applications in software engineering. For instance, Gao et al. (Gao et al., 2023) investigated the influence of the selection, sequence, and quantity of demonstration examples on in-context learning performance in code intelligence tasks, namely, code summarization, bug fixing, and program synthesis. They found that meticulously curated demonstrations can significantly outperform conventional demonstration construction techniques. Different from the aforementioned research, we focus on designing a new prompt engineering method to synthesize compilable code for incomplete code snippets on Q&A sites by leveraging LLM as a knowledge base. We introduce ZS4C, a lightweight zero-shot approach that decomposes the task into two stages, and enables a collaborative conversation between the LLM and a compiler to resolve compilation errors.
3. Methodology
Figure 1. Overview of our approach ZS4C.
We present the overview of our approach ZS4C in Figure 1. ZS4C is comprised of two components: Import Statement Inference and Conversational Error Fixing. In Import Statement Inference, we leverage a large language model (LLM) to identify the missing dependent libraries (by identifying missing import statements) in the code snippet. In Conversational Error Fixing, we further fix compilation errors caused by the wrong dependent libraries inferred by Import Statement Inference and other compilation errors (e.g., syntax error) to boost the compilation success rate of a code snippet, through the conversation between a compiler and the LLM. For simplicity, in the following section, we refer to Import Statement Inference as ImportStInfer and Conversational Error Fixing as ConversationFixing.
3.1. Import Statement Inference
Previous studies report that the majority issue that prevents code snippets on Q&A sites from compiling successfully is the absence of dependent libraries (Terragni et al., 2016; Dong et al., 2022). Therefore, our first step is to infer the missing library import statements for a given code snippet. At a high level, ImportStInfer takes a code snippet $input\_code\_snippet$ as input and outputs the missing import statements for $input\_code\_snippet$. More specially, as shown in Figure 1, we first construct a prompt by using our designed simple yet effective prompt template. We then apply the self-consistency method (Wang et al., 2022) to obtain more reliable and coherent results from LLM. We elaborate on each step below.
3.1.1. Prompt design
Prompts are user-provided inputs, such as queries and instructions that guide LLMs and instruct their behavior for specific tasks. Previous studies have demonstrated that the quality of input prompts plays a crucial role in the performance of LLMs, significantly influencing the quality of the output (Gao et al., 2021; White et al., 2023; Zhou et al., 2022b; Min et al., 2022). A prompt template is often used to generate prompts in a consistent and reproducible way. We follow the style of prior work (Li et al., 2023; Xia and Zhang, 2023) in providing a short and concise prompt template tailored to the import statement inference task, as illustrated below.
#Instruction Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Do not check for any import statements in the code. Only give correct imports by not using wildcard imports. Please note that you need to pay close attention and your response should be specific and accurate. Avoid repetition and must not generate wrong and nonexistent imports:” #Query {$input\_code\_snippet$}
In Instruction, we first use Presentation_Guide “Reply with to-the-point answer, no elaboration” to guide the model to provide a compact and to-the-point answer to our query. Without this directive, the model occasionally includes explanations in the output, which brings additional workload for extracting the missing import statements from the output.
We design the instructions to guide and control the output of the model for our task step by step in Main_Instruction. We first ask the model not to check the code for any import statements if present. [^2] Note that, in line with previous studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), the input code snippets utilized for evaluation in this study do not contain any import statements. This design directs the model’s attention exclusively to the body of the code rather than any existing import statements, which could be incorrect. Moreover, we observed that the model sometimes generates responses including “no imports are found”, likely stemming from a misinterpretation of the task. To mitigate this issue, we explicitly instruct the model to avoid seeking import statements from the input code snippet. Next, we instruct the model to provide only correct imports and to avoid using wildcard imports. Wildcard import refers to the mechanism by which one can import all classes and interfaces from a particular package without specifying their fully qualified names by using the asterisk (*) character. For instance, in Java, one can use “import java.util.*;” to import all classes from the “java.util” package, including the ArrayList, HashMap, and Scanner classes, among others, rather than individually importing classes, e.g., “import java.util.ArrayList”. We instruct the model to avoid wildcard imports for several reasons: 1) wildcard imports may cause naming conflicts if multiple classes with the same name exist in different packages; 2) we want to steer the model towards identifying the exact classes required rather than resorting to the short-cut wildcard imports; 3) wildcard imports are generally discouraged in practice because it can compromise the code readability and introduce unintended side-effects that can be very difficult to debug (Van Oort et al., 2021). Following the recommended best practice for prompt design mentioned in  (White et al., 2023), we add a sentence “Please note that you need to pay close attention and your response should be specific and accurate.” to explicitly emphasize the importance of specificity and accuracy for the LLM to prioritize. We also observed that in some cases, the model generates duplicate import statements and incorrect import statements that referenced non-existent libraries. To mitigate this issue, we added a sentence “Avoid repetition and must not generate wrong and nonexistent imports” to the prompt template.
Finally, we instantiate the {$input\_code\_snippet$} with the actual code snippet and send the filled prompt to LLM for inference.
3.1.2. Self-consistency
LLMs are nondeterministic by nature and would respond differently to the same input prompts (Wang et al., 2022; OpenAI, 2023b). Many LLMs, including ChatGPT, provide a temperature hyper-parameter, which ranges from 0 to 1, to control the creativity or randomness of the generated response. A higher temperature (e.g., 0.7) results in more diverse and creative output while a lower temperature (e.g., 0.2) makes the output mode deterministic and focused. However, we observed that, even with a low temperature of 0, the import statements generated by LLM could vary time by time with the same prompt. For instance, using the same prompt for a Java class we got three different outputs: 1) “The code does not require any additional import statements.” 2) “The code requires the following imports: java.text.NumberFormat; java.util.Locale”. 3) “import java.text.NumberFormat;
import java.util.Locale;”.
To obtain reliable and coherent results from the model, we employ the self-consistency method proposed by Wang et al. (Wang et al., 2022). As indicated by its name, self-consistency simply asks a model to run the same prompt multiple times and determines the final answer based on the principle of frequency majority. More specifically, we run the same prompt for $K$ times and the model generates $K$ lists of import statements for our query. We count the frequency of each list and take the list with the highest frequency as the final result. If at least two lists have the same frequency, we repeat the process (i.e., running the same prompt for another $K$ times) until we break the tie. For instance, we query the model using the same prompt 5 times, and the model generates five lists, $L_{1}=\{A,B,C\}$, $L_{2}=\{B,A,C\}$, $L_{3}=\{A,B,C,D\}$, $L_{4}=\{A,B\}$, and $L_{5}=\{C,B,A\}$, we take $\{A,B,C\}$ as the final result as it (the combination of three libraries - A, B, C) appears 3 times with the highest frequency. In this study, we empirically set $K$ to 10, which achieves the best performance (see results of RQ4).
3.2. Conversational Error Fixing
Through ImportStInfer, we could identify potential missing import statements. However, the code may still not be compiled due to incorrect import statements, bad syntax, typos, and undeclared variables. Therefore we need further validation of our generated code and boost the compilation rate by fixing those errors. Drawing inspiration from recent advancements in the Multi-Agent Conversation and its applications across various tasks (Wu et al., 2023; Zhang et al., 2023; Xi et al., 2023), where multiple LLM-based agents work collaboratively and outperform a single LLM-based agent, we design a ConversationFixing component to further improve the results given by ImportStInfer. Within this component, a Java compiler is integrated into workflow with the LLM (via conversations) to address compilation errors. This compiler not only validates the generated code but also produces an error log that aids in resolving compilation issues. Specifically, we leverage the feedback, i.e., the error log, from the compiler’s validation as context, prompting the LLM to fix the errors for the given code snippet.
Input : Code snippet with inferred import statements, $CS$
Input : Maximum number of interaction between compiler and LLM, $M$
Input : Large language model, $LLM$
Output : Most recent fixed Code, $FC$
1 $errorLog$ = ””
2 $history$ = ””
3 // Conversation between compiler and LLM up to $K$ times
4 for $i\leftarrow 0$ to $M-1$ do
5       // load dependent libraries dynamically
6       Compiler.loadExternalLibraries($CS$)
7       $CompileResults$ = Compiler.Compile($CS$)
8       // $CS$ is failed to compile
9       if $CompileResults$ is failed then
10             $errorLog$ = getErrorLog(CompilerResults)
11             $Prompt$ = constructPrompt($errorLog$, $history$)
12             $CS$, $history$ = queryLLM($LLM$, $Prompt$)
13
14      else
15             break
16       end if
17
18 end for
[…]
return $FC$
Algorithm 1 Algorithm for Conversational Error Fixing
We present the logic of ConversationFixing in Algorithm 1. The algorithm takes the code snippet with inferred import statements generated by ImportStInfer, denoted as $CS$, the maximum number of interactions between compiler and LLM, represented by $M$, and a large language model, $LLM$, as input, and outputs a most recent fixed code, denoted as $FC$. First, we load external libraries and compile the $CS$ (lines 6-7) to examine if $CS$ can be compiled successfully. If the compilation is successful, we return $CS$ directly. If not, we construct a prompt incorporating the error log and $CS$ (lines 10 - 11). Subsequently, We query the $LLM$ with the constructed prompt, seeking corrections for $CS$ (line 12). This conversation (lines 5 - 15) is repeated up to $M$ times until a successful compilation.
Note that during the compilation, the compiler needs to search for the external libraries to import as external dependencies (line 6). To do so, we dynamically import the external libraries based on the import statements obtained from ImportStInfer or previous round of ConversationFixing. More specially, we construct a knowledge base to store the mappings between fully qualified names (FQN) of classes/functions and their corresponding external libraries (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar) in an offline manner. See more details of knowledge base construction in Section 4.2. When a piece of code with inferred import statements is input into ConversationFixing, we search the external libraries for each FQN from the import statements and load the corresponding libraries from the constructed knowledge base. It’s worth noting that a single FQN possibly could correspond to multiple external libraries (e.g., different versions of the same libraries provide the same class) (Dong et al., 2022; Subramanian et al., 2014). To mitigate such a case, we selected the one that satisfies pre-defined constraints following the previous study (Subramanian et al., 2014). In our study, we consider the invoked functions and attributes as the constraints. If there exist multiple libraries available after resolving constraints, we select the one that was released most recently. We assume it is realistic for a user to use the most recent libraries. Note that our approach does not guarantee to find all external libraries due to various reasons (e.g., incorrect FQN in import statements), instead our goal is to leverage the message from the compiler to help LLM infer the correct import statements iteratively. For instance, the example in Figure 2 (left), for the import statement “import org.joda.time.PeriodFormatterBuilder;”, no external library was found since the import org.joda.time.PeriodFormatterBuilder (missing format) is incorrect. We still compile the code and the error message generated by the compiler helps LLM fix the error (see more detailed results in Section 5.2).
Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Now fix the error by focusing on fixing the import statements by not using wildcard imports and must not modify code body which means do not change anything inside the class. So, it can be successfully compiled and reply with full code.” # Prompt template for the 1st attempt: {Presentation_Guide} “See the code below:” $\{input\_code\_snippet\_with\_inferred\_import\_statements\}$ “For the above code I got the below error log:” $\{error\_log\}$Main_Instruction} # Prompt starting from the 2nd attempt: $Context$: $\{chat\_history\}$$\{last\_code\_from\_ChatGPT\}$ “You gave the above imports fix in your attempt $\{attempt\_number\}$. But compiler gave this error:” $\{last\_error\_message\}$Main_Instruction}
In each iteration, we update the $CS$ as suggested by $LLM$ and provide the chat history as contextual information. If, after attempting $M$ times, $CS$ remains non-compilable, the latest revised code is returned. Our design mimics how developers use an LLM (e.g., ChatGPT) to fix compilation errors for a piece of code over conversation with LLM (e.g., ChatGPT).
We present our designed prompt template-2. We designed two templates: one for the first attempt and the other for subsequent attempt(s). In the first attempt, we attach the uncompilable code with an error log and ask LLM to fix the error and return a fixed version. Similar to the Prompt Template - 1 used in ImportStInfer (see Section 3.1), we ask LLM to avoid wildcard imports. We observed cases where the LLM circumvents errors by simply removing code segments that cause the compilation error. To mitigate this issue, we instruct the LLM to not modify the code body. For the following attempt, we attach the history from the previous attempts and ask LLM to fix based on history. Since the size of LLM’s context window is limited, we only keep the most recent chat history to offer the context that fits the window size.
4. Experimental Design
In this section, we present research questions (RQs), dataset, evaluation metrics, our analysis approach for RQs, and implementation details.
4.1. Research questions
We evaluate ZS4C in different aspects to answer the following research questions.
  • RQ1: How effective is our approach ZS4C?
  • RQ2: What is the contribution of individual components to ZS4C?
  • RQ3: Does few-shot learning improve ZS4C?
  • RQ4: How do parameters affect the effectiveness of ZS4C?
4.2. Datasets
We use an established benchmark called StatType-SO, which was used in the related studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019). The dataset consists of 267 code snippets of six popular Java libraries. Table 1 summarizes the basic statistics of StatType-SO. The six libraries are of different sizes and cover various application domains. This dataset was manually collected from Stack Overflow posts and represents the scenario in which developers use a wide variety of real Java libraries in practice (Phan et al., 2018). This benchmark consists of the code snippets and their dependent libraries. For constructing the knowledge base that stores the mapping between fully qualified names (FQN) of classes/functions and their corresponding external libraries, we downloaded all versions of the most popular 500 libraries from maven repositories [^3] https://mvnrepository.com/popular. We extracted the FQN for all classes/functions from the jar files and created the knowledge base.
Table 1. Basic statistics of StatType-SO.
#Code Snippet #Average LOC #Unique Import Statements
Android 50 38 129
JDK 23 90 103
Hibernate 5 50 123
JodaTime 50 19 48
GWT 50 35 108
XStream 44 37 80
4.3. Evaluation metrics
To evaluate the effectiveness of ZS4C as compared with alternatives, we consider the following two criteria.
Criterion 1: Compilation Rate on the Synthesized Code. To measure the effectiveness of ZS4C in synthesizing compilable code from incomplete code snippets, we use compilation rate, i.e., the ratio of code snippets that successfully compile post-repair to the total number of snippets evaluated, following previous studies (Dong et al., 2022; Terragni et al., 2016).
Criterion 2: Accuracy of the Synthesized Import Statements. A trivial fix, such as removing dependent libraries and related code can also resolve compilation errors. However this would risk significantly altering the code’s semantics. Thus, in addition to the compilation rate, we assess the accuracy of the synthesized import statement list by comparing them to the ground truth list. To qualify the alignment between the two import statement lists, we define metrics True Positive (TP), False Positive (FP), False Negative (FN) as:
  • True Positive (TP): Denotes the count of import statements in the ground truth that are correctly predicted in the synthesized list.
  • False Positive (FP): Denotes the count of import statements in the synthesized list that are not present in the ground truth.
  • False Negative (FN): Denotes the count of import statements missing in the synthesized list but present in the ground truth.
[…]
Following the baseline SnR (Dong et al., 2022), we also measure a list-wise matching degree for each library. Specifically, for each code snippet, the synthesized result is classified into one of the following categories:
  • Same: The approach synthesizes the exact same expected set of import statements.
  • Different: The approach synthesizes one or more alternatives for some expected import statements.
  • Missing: The approach synthesizes a subset of the expected set of import statements.
  • Extra: The approach synthesizes a superset of the expected set of import statements.
  • None: The approach synthesizes no import statements.
Finally, we summarize the distribution of the above categories for each testing library.
4.4. Base large language model
In this study, we choose ChatGPT (OpenAI, 2023a) as our base LLM. We select ChatGPT as it is recognized to be one of the most powerful LLMs and has gained wide attention in recent months by demonstrating the outstanding capability of solving tasks in various areas (Frieder et al., 2023), including software engineering (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023; Xia and Zhang, 2023). We accessed ChatGPT, i.e., the default GPT-3.5-Turbo model, via the OpenAI API (OpenAI, 2023b).
4.5. Implementation details
Experiments were conducted on a Linux server equipped with one Nvidia RTX 3090 GPU, a 24-Core CPU, and 64 GB of RAM. The implemented code has been made available on GitHub. We use the default setting (i.e., Maximum length = 256, Top P = 1, Frequency penalty = 0, and Presence penalty = 0) for ChatGPT except for temperature. For self-consistency, we set a high temperature of 1 following the recommendation by previous study (Wang et al., 2022). For other components that involve ChatGPT, we set a medium temperature of 0.5.
4.6. Approach of RQs
4.6.1. Approach of RQ1
We compare ZS4C with the following two baselines:
  • SnR: The state-of-the-art technique for synthesizing incomplete code snippets by inferring missing import statements and making them compilable (Dong et al., 2022). SnR has been demonstrated to archive better performance than other approaches (e.g., COSTER (Saifullah et al., 2019), Baker (Subramanian et al., 2014), and StatType (Phan et al., 2018)) on StatType-SO. Therefore, in this study, we only compare with SnR. We leveraged the replication package of SnR as provided in their paper. [^4] We found disparity between the number reported in the paper and the number derived from their replication package. We present the numbers derived from their replication package in this paper. We do not compare with CSNIPPEx because this Eclipse plugin is no longer accessible.
  • BasePrompt: We also consider a baseline that employs a straightforward prompt to instruct a LLM to make the given code snippet compilable. Specifically, we replace the Main_Instruction as “Make the code below compilable: {$input\_code\_snippet$}” in Prompt Template-1 (ref. Section 3.1) and while retaining the remainder of the prompt template unchanged.
We evaluate the effectiveness of ZS4C against the two baselines on the two criteria outlined in Section 4.3. To ensure the reliability of our results and mitigate potential biases arising from the inherent randomness of the LLM, we run both ZS4C and BasePrompt five times and take the median value as the final result.
4.6.2. Approach of RQ2
[…]
4.6.3. Approach of RQ3
Few-shot in-context learning is a paradigm where a limited set of examples are provided in the prompt and allows an LLM to learn from the examples in a specific context. This method has demonstrated its effectiveness in enhancing LLM performance, especially when compared to scenarios where no examples (zero-shot) are provided for various downstream tasks (Nashid et al., 2023; Brown et al., 2020; Gao et al., 2021; Min et al., 2022). Therefore, in this RQ, we aim to investigate if enabling this option would benefit ZS4C.
Selecting good demonstrations is essential for few-shot in-context learning. Prior studies show that selecting demonstration examples with higher similarity (to the input) can improve in-context learning performance (Nashid et al., 2023; Zhou et al., 2022a). Thus, in this RQ, we incorporate code snippets that are most similar to each input code snippet, along with their corresponding import statements, as demonstrations within the prompt. Specifically, we employ the commonly used CodeBERT (Feng et al., 2020) model that was trained for code-related tasks to learn a vector representation for each code snippet in StatType-SO. Based on the code representations output by CodeBERT, we select the code snippets presenting the highest cosine similarities with the input code snippet. Due to the window size token limitation of ChatGPT (i.e., 4k), we only put the most similar code snippet and its import statements in the prompt as the demonstration example. We put the retrieved demonstrated example after the Main_Instruction in Prompt Template - 1 as shown in Section 3.1.1. In other words, we create a one-shot version of ZS4C.
4.6.4. Approach of RQ4
ZS4C has two parameters: $K$ the number of sampling in self-consistency, and $M$ the maximum rounds of conversation between the compiler and the LLM in Conversational Error Fixing. In this RQ, we investigate the impact of each parameter on the effectiveness of ZS4C, where we set $K$ to 10, 20, and 30 and set $M$ to 5, 10, and 15, respectively.
5. Results
5.1. Results of RQ1
Table 2. The comparison of SnR, BasePrompt, and ZS4C in terms of F1, precision (Pre), and recall (Rec), and compilation rate (CR). We present the average for F1, Rec, Pre, and the total compilation rate in the summary row. The cells that have the best performance are marked in bold.
SnR BasePrompt ZS4C
F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
Android
0.952
0.909
1
37
0.967
0.95
0.987
44
0.990
0.990
0.990
48
JDK
0.887
0.796
1
12
0.482
0.317
1
15
0.997
1.000
0.994
23
Hibernate
0.948
0.918
0.980
23
0.75
0.606
0.985
22
0.902
0.834
0.982
32
Joda-Time
0.860
0.759
0.992
37
0.979
0.971
0.988
44
0.985
0.994
0.977
50
GWT
0.835
0.721
0.991
30
0.951
0.927
0.977
32
0.968
0.940
0.997
46
XStream
0.934
0.877
1
29
0.863
0.78
0.967
22
0.934
0.890
0.980
35
Summary
0.903
0.830
0.994
168 (62.9%)
0.832
0.759
0.984
179 (67%)
0.963 0.938
0.987
234 (87.6%)
[…]
Same Different Extra Missing None
SnR 143 78 4 35 7
BasePrompt 173 55 14 19 6
ZS4C 202 35 16 14 0
ImportStInfer 191 48 14 12 2
[…]
ConversationFixing 175 48 22 19 3
We present the comparison results among ZS4C and two baseline approaches in Tables 2 and  3.
The experimental results show that ZS4C successfully synthesized 234 compilable code snippets out of the 267 provided incomplete code snippets, resulting in an impressive overall compilation rate of 87.6%, outperforming two baselines. Compared with SnR, ZS4C showed a 39.3% improvement in terms of compilation rate. [^5] SnR paper reports that they can compile 197 out of 267 (73.8%) code snippets successfully after the repair. However, in their replication package, we found 168 code snippets were marked as comparable. Nevertheless, our approach outperforms SnR by improving the number of compilable code snippets to 234. Furthermore, ZS4C consistently outperformed the two approaches for all libraries. A major factor behind this boosting in compilation rate is the ConversationFixing component, which enables conversations between the compiler and the LLM. A deeper dive into the effectiveness and impact of this component can be found in RQ2 and Section 6.1.
When measuring the accuracy of the synthesized import statements, on average, ZS4C achieves 0.961, 0.987, and 0.938 in terms of F1, Precision, and Recall, respectively, as detailed in Table 2. Compared with SnR, ZS4C achieves an improvement of 6.6% and 13.4% in terms of F1 and precision. In terms of Recall, ZS4C has a slightly lower performance than SnR, which achieves an almost perfect precision of 0.994 due to its constraint-based inference. In six of the target libraries, ZS4C outperforms SnR in five, except Hibernate. We perform failure cases analysis to understand why eventually some cases are not compilable after the fixing (see more details in Section 6.1).
Compared with BasePrompt, ZS4C can compile 55 more code snippets, reflecting a 30.7% enhancement in the compilation rate. Furthermore, ZS4C elevates the accuracy of import statement inference, with the F1 score rising from 0.832 to 0.961. These improvements indicate the significance of task-specific prompt engineering in optimizing the performance of LLMs like ChatGPT.
To complement accuracy metrics, we also compare the list-wise matching degree. ZS4C (202) can infer more exact match import statement lists (same) than SnR (143) [^6] SnR paper reports that they can achieve 183 same, 62 different, 11 extra, 4 missing, and 7 none matches. However, in their replication package, we found the number we reported in the table. Nevertheless ZS4C can achieve more exact same matches. and BasePrompt (193). Also ZS4C can always output thing that is related to the correct import statements (none is 0).
To understand why ZS4C outperforms SnR, we manually study the cases where ZS4C performs better than SnR. We observe two reasons. First, SnR needs to repair an input code snippet to a syntactically valid compilation unit and then generate the Abstract Syntax Tree (AST) for the following procedures. However, the syntax repairing is not 100% guaranteed. Therefore, in certain cases, SnR failed in the first repairing step. While ZS4C does not rely on static analysis (e.g., parsing AST). LLM directly takes code snippets as sequence input and does the inference, which reduces the risk of failures due to errors incurring in static analysis. Second, SnR infers the appreciated import statements by solving constraints. However, if multiple libraries meet the constraints simultaneously, SnR does not have a strategy to handle this. While ZS4C leverages LLMs as the knowledge base to perform the inference, LLMs have sufficient knowledge to decide which libraries are most appreciated given the context.
ZS4C improves the compilation rate of the SOTA approach, SnR, from 63% to 87.6%. On average, ZS4C’s synthesized import statements are more accurate than SnR, with an improvement of 6.6% in the F1 score. Furthermore, ZS4C is more effective than BasePrompt, which represents a straightforward way of utilizing ChatGPT for this task.
5.2. RQ2: What is the contribution of individual components to ZS4C?
[…]
F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
Android
0.948
0.915
0.982
39
0.980
0.980
0.980
42
0.927
0.902
0.954
43
JDK
0.919
0.850
1
19
0.997
1
0.994
23
0.825
0.707
0.992
21
Hibernate
0.886
0.813
0.974
27
0.898
0.828
0.981
38
0.876
0.803
0.963
34
Joda-Time
0.964
0.941
0.988
42
0.967
0.959
0.976
29
0.968
0.965
0.970
49
GWT
0.939
0.905
0.976
35
0.956
0.924
0.990
45
0.888
0.879
0.896
44
XStream
0.884
0.936
0.837
31
0.910
0.850
0.980
28
0.898
0.850
0.951
34
Summary
0.923
0.893
0.960
193 (72.2%)
0.951
0.924
0.984
205 (76.8%)
0.897
0.851
0.954
225 (84.2%)
[…]
We additionally observe that ConversationFixing achieves a higher compilation rate (84.2%) than ImportStInfer (72.2%). This underscores the benefits derived from the collaborative conversations between the LLM and the compiler. Indeed, ConversationFixing simulates a practical scenario where a developer interacts with ChatGPT through several rounds to address the compilation errors in their code. Yet, when we shift the focus to the efficacy of both components in inferring missing import statements, the tables turn. ImportStInfer has a F1 of 0.951, which is much higher than ConversationFixing’s F1 score of 0.897. This demonstrates that extracting all required important statements from the code body can be difficult when the primary goal of the LLM is to resolve compilation errors. By breaking the task into two steps—first inferring import statements and then iteratively fixing compilation errors—the efficiency of the LLM is enhanced. As shown in Table 4, when isolated, neither component approaches the performance of ZS4C in terms of the two considered criteria. This shows that both components significantly contribute to ZS4C, and their combined strengths boost ZS4C’s performance.
To further understand the effectiveness of the ConversationFixing component after obtaining the inferred import statements, we compare ImportStInfer and ZS4C. We observe that the two approaches exhibit a comparable proficiency in inferring missing import statements, i.e., their F1 scores are 0.953 and 0.964, respectively. However, ZS4C achieves a higher compilation rate (87.6%) compared with ImportStInfer (72.2%) with an improvement of 21.3%. This observation confirms the effectiveness of the ConversationFixing component in fixing compilation issues and thus boost the ability of ZS4C in synthesizing compilable code significantly.
To gain insight into the corrections made by the ConversationFixing component, we conducted a qualitative analysis, comparing code snippets before and after processing through this component. Specifically, we manually reviewed all 29 cases where the compilation errors were resolved by the ConversationFixing component. We observe that in 26 cases, ConversationFixing either added the missing imports or corrected the wrong imports. For instance, Figure 2 presents a code snippet that is fixed by ConversationFixing. After going through the Import Statement Inference, four import statements are inferred. However, after being validated by the compiler, it was reported that two of imported libraries, namely org.joda.time.PeriodBuilder and org.joda.time.PeriodFormatterBuilder, were not recognized. The error log is then sent back to ConversationFixing and eventually the compilation issue is fixed by removing the two incorrect imports and add a correct one. In the remaining three cases, we observe that LLM modifies the code body and plays tricks to make the code compilable although we explicitly instruct the model not to change code body in our prompt template. For instance, in a code snippet utilizing the GWT library, there is a variable $table$ whose type is Table, LLM removed $table$ from the code and made the code compilable.
Figure 2. An example of code snippet that is fixed by ConversationFixing. The left example is before and the right example is after applying the ConversationFixing.
Both components in ZS4C contribute significantly to its performance. For instance, by employing the Conversational Error Fixing component, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement. Additionally, the self-consistency method improves the effectiveness of the Import Statement Inference component.
5.3. RQ3: Does few-shot learning improve ZS4C?
Table 5 presents the results of ZS4C in two settings, zero-shot and one-shot. Surprisingly, one-shot in-context learning does not outperform zero-shot in any evaluation metrics. Instead, there is a slight decline in terms of F1. We suspect the reason could have originated from the advantages of using demonstration examples. Demonstration examples are usually provided for in-context learning for two purposes: 1) to help LLM understand the intent of a task and learn the patterns to complete the task (White et al., 2023; Webson and Pavlick, 2021; Lu et al., 2021); 2) to provide domain-specific knowledge for downstream tasks (Nashid et al., 2023; Gao et al., 2023; Zhou et al., 2022a). In our cases, apparently, none of the above advantages of providing examples hold. Our task is straightforward, and the model can understand the intent expressed in the zero-shot prompt precisely. This is possible because ChatGPT is trained on a large dataset Common Crawl, which includes data from Stack Overflow and Github (Nagel, 2022). ChatGPT could understand the task precisely. Meanwhile, the demonstrations do not reveal any domain-specific knowledge that can be utilized.
Table 5. The comparison of the performance of ZS4C in one-shot and zero-shot settings.
Compilation Rate F1 Recall Precision
one-shot 86.9% 0.955 0.934 0.981
zero-shot 87.6% 0.961 0.938 0.987
Surprisingly, one-shot in-context learning does not improve the performance of ZS4C.
5.4. RQ4: How do parameters affect the effectiveness of ZS4C?
Figure 3. The impact of parameters $K$ (left) and $M$ (right) on ZS4C.
Figure 3 (left) presents the results of parameter sensitivity analysis on $K$ (i.e., number of sampling in self-consistency). As $K$ increases, the effectiveness of ZS4C to infer the import statements declines slightly, from 0.961 to 0.934 in terms of F1. We can observe similar trends for accuracy, recall, and precision. However, we observe the compilation rate improves slightly from 0.873 to 0.891. One possible explanation is that ConversationFixing provides a safe net for ImportStInfer even if it makes mistakes for some cases.
When we look at $M$ (i.e., the maximum round of conversation between LLM and compiler) as shown in Figure 3, we notice that the effectiveness of inferring import statements remains stable when $M$ increases from 5 to 15, i.e., the F1 value ranges from 0.962 to 0.966. While, the compilation rate increases from 0.873 to 0.902 when $M$ increases from 5 to 10, and then remains stable. This indicates in some cases, conducting more rounds helps with issue fixing.
We then delved deeper to understand how compilation errors are addressed over multiple rounds of conversations in Conversational Error Fixing. We observe that more than half of the cases (15/29) could be fixed in the first round. However, in some cases, even more rounds of conversation are conducted, Conversational Error Fixing eventually fails to complete the task. For instance, in one code snippet of the Hibernate project, as shown in Figure 4, there is a private variable called userGroup with type $UserGroup$. After the first round of conversation, all the libraries except the one for $UserGroup$ are identified (the left synthesized code snippet). Therefore, in the second attempt, Conversational Error Fixing asked LLM to fix the issue for the not recognized type $UserGroup$. As shown in Figure 4 (right), LLM added a statement to import hibernate.UserGroup. However, hibernate.UserGroup is not a real class that exists in the hibernate library, which is fabricated by LLM. In the following rounds of conversation, LLM kept in a loop of removing and adding this import statement again and again.
Figure 4. An example of eventually failed after 15 rounds of conversations.
Overall, the effectiveness of ZS4C is not sensitive to its parameters.
6. Discussion
6.1. Failure cases analysis
To further understand the reasons that contribute to the failure of ZS4C in synthesizing compilable code even after applying Conversational Error Fixing, we conduct a qualitative analysis of the failed cases. We collect all 33 failed cases and manually examine the conversation history between the compiler and LLM. We derived the following five reasons.
  • Unconstrained class (14): The LLM fails to infer the classes because such classes have little/no constraints to follow except the name. For instance, the type $UserGroup$ in the example of Figure 4 does not have any constraints, such as fields and methods. Consequently, any library could be imported, as long as it has a class called $UserGroup$.
  • Partial inference (13): The LLM fails to infer the exact same FQN (fully qualified name) of classes. This usually happens when a class has a long FQN of a class. For instance, in a case in Hibernate, the correct library to import is com.thoughtworks.xstream.converters.basic.AbstractSingleValueConverter, which contains 73 characters. The LLM synthesized com.thoughtworks.xstream.converters.enums.AbstractSingleValueConverter.
  • Fake inference (3): The LLM produces a fake library derived from the given code context. For instance, in a case of Android, it passes $R.id.lay$ as a parameter to a locally defined function $LinearLayoutfindViewById(intlay){}$. Based on the context, the correct FQN for $R.id$ is $Android.R.id$. However, the LLM generated a fake one called $AndriodExamples.R$, with $AndriodExamples$ being a package name present in the code snippet.
  • Alternative inference (2): The LLM generates an incorrect yet real library to import.
  • Unexpected code modification(1) : The LLM alters the original code body, leading to unintended deviations.
We observe 14 cases (42.4%) failed due to unconstrained class and 11 of them are in Hibernate. This is reasonable because the code snippets of Hibernate in StatType-SO usually have unconstrained classes (e.g., user-defined classes), which only have a class name defined in the code without any constraints on its field and method. The inability to manage unconstrained classes is also a prevalent issue in previous methodologies, including constraints-based approaches (Dong et al., 2022; Subramanian et al., 2014) and statistical model-based approaches (Phan et al., 2018; Saifullah et al., 2019). A simple way to mitigate this issue is to create an empty class for such a class. We instruct the LLM to fix compilation errors without altering the original code body, aiming to avoid trivial fixes. Consequently, ZS4C is unable to create empty classes.
We also observe that 13 cases (39.3%) failed due to partial inference. This is a common issue for all generative models (Huang et al., 2022). The probability of accurately predicting a specific sequence of tokens diminishes as the sequence lengthens. Most of the partial inference happens in GWT and XStream and we found the ground truth import statements in those two libraries are longer than the other libraries. Specifically, code snippets from GWT and XStream have an average of 46 and 48 characters, respectively. The code snippets from the other four libraries typically range between 29 and 35 characters. Although, our ZS4C cannot infer the exact correct import statements for a few cases and eventually leads to failure, the inferred import statements can be used to identify a corresponding existing library by employing keyword searches or by calculating similarity scores.
Interestingly, we observed one case, where the original code body was modified by LLM in the second round of conversation in ConversationFixing, and the fixing process turned in the wrong direction because the rest of the fixing was based on the modified code. Specifically, in a GWT code snippet shown in Figure 5, a class named $LayoutContainer$ was changed to $Container$ in the 2nd round of fixing. This class was further changed to $SimpleContainer$ in the 3rd round. The fixing process turned in the wrong direction after the 1st round even if all the correct import statements could be inferred and eventually get compilable.
Figure 5. An example of the original code body is modified by LLM.
6.2. Limitations of ZS4C and potential solutions
In this section, we discuss the limitations of our approach and the potential solutions for mitigating that limitations in the future.
Hallucination. Both the fake inference and partial inference, discussed in Section 6.1, can be attributed to a prevalent issue with LLM - hallucination, where the model speaks false knowledge as if it is accurate (Bang et al., 2023). ZS4C is grounded on LLM, naturally, it inherits this problem. A potential solution to this challenge is the use of retrieval-augmented generation (RAG) (Lewis et al., 2021), which retrieves facts from an external knowledge base to ground LLM on the most accurate, up-to-date information. In the context of our task, the supplementary knowledge could be sourced from API documentation relevant to the code snippet.
Unexpected code modification from LLM. We observed that in four cases, ZS4C modified the original code during the conversation between the compiler and LLM although we already emphasized in the prompt to instruct ChatGPT to keep the original code. This is due to the stochastic nature of LLM, i.e., the output cannot be totally guaranteed by the prompt (Brown et al., 2020). One possible solution to this problem is to validate if the code snippet generated by LLM is still the same as the original code. We encourage future research to investigate this direction.
6.3. Threats to validity
Threats to external validity. Consistent with prior studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), we utilize the StatType-SO benchmark as the evaluation dataset. This dataset, comprising manually verified code snippets from six prominent Java libraries across diverse domains, may not be representative of libraries in all programming languages. Nonetheless, ZS4C is language-agnostic and it does not rely on Java-specific constructs. We have chosen ChatGPT (GPT-3.5-Turbo) as the base LLM for our study. Consequently, our findings might differ if alternative LLMs, such as GPT-4, are employed.
Threats to internal validity. The primary threat to internal validity comes from the manually crafted prompt templates tailored for the task in ZS4C. We cannot guarantee the optimality of our prompts, nor can we consider all the potential prompts. To mitigate this threat, we elaborate on all of our design concerns and compare ZS4C with multiple baselines when basic prompts are utilized. Given the context window token constraints of ChatGPT, we consider adding up to one demonstration example to ZS4C and form its one-shot version, as incorporating more examples might breach the token limit. Furthermore, only the most recent chat history is retained in ConversationFixing once the context window limit is approached.
7. Conclusion
We proposed ZS4C, a lightweight approach that can synthesize compilable code for incomplete code snippets on Q&A sites leveraging an LLM. ZS4C consists of two components, one for automatic inference of import statements and the other for further fixing import statements and other compilation errors via iterative conversation between the LLM and a compiler. We have carefully designed task-specific prompts for the two components and applied the widely adopted self-consistency method to improve the robustness of the inference components.
We compared ZS4C with a state-of-the-art approach SnR on a benchmark dataset. Our experimental results show that ZS4C improves the compilation rate of the SOTA approach, SnR, from 63% to 87.6%. We conducted an ablation study to understand the contribution of each component in ZS4C and investigated the impact of few-shot in-context learning. The ablation study revealed that both components in ZS4C play crucial roles in the effectiveness of ZS4C. We also conducted a qualitative study to analyze the failure cases and point out future directions to improve ZS4C.
8. Data Availability
We made our replication package publicly available (dat, 2023) to encourage future research on improving the synthesis of compilable code from incomplete code snippets.

### v3

@@
-ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT
+MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=
 Abstract.
-Technical question and answering (Q&A) sites such as Stack Overflow have become an important source for software developers to seek knowledge.
-However, code snippets on Q&A sites are usually uncompilable and semantically incomplete for compilation due to unresolved types and missing dependent libraries, which raises the obstacle for users to reuse or analyze Q&A code snippets.
-Prior approaches either are not designed for synthesizing compilable code or suffer from a low compilation success rate.
-To address this problem, we propose ZS4C, a lightweight approach to perform zero-shot synthesis of compilable code from incomplete code snippets using Large Language Model (LLM).
-ZS4C operates in two stages.
-In the first stage, ZS4C utilizes an LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template.
-In the second stage, ZS4C fixes compilation errors caused by incorrect import statements and syntax errors through collaborative work between ChatGPT and a compiler.
-We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO against the SOTA approach SnR.
-Compared with SnR, ZS4C improves the compilation rate from 63% to 87.6%, with a 39.3% improvement.
-On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1.
-imcomplete code snippets, Large Language Model, Program synthesis, API inference, Prompt Engineering, ChatGPT
+Technical Q&A sites are valuable for software developers seeking knowledge, but the code snippets they provide are often uncompilable and incomplete due to unresolved types and missing libraries.
+This poses a challenge for users who wish to reuse or analyze these snippets.
+Existing methods either do not focus on creating compilable code or have low success rates.
+To address this, we propose ZS4C, a lightweight approach for zero-shot synthesis of compilable code from incomplete snippets using Large Language Models (LLMs).
+ZS4C operates in two stages: first, it uses an LLM, like GPT-3.5, to identify missing import statements in a snippet; second, it collaborates with a validator (e.g., compiler) to fix compilation errors caused by incorrect imports and syntax issues.
+We evaluated ZS4C on the StatType-SO benchmark and a new dataset, Python-SO, which includes 539 Python snippets from Stack Overflow across the 20 most popular Python libraries.
+ZS4C significantly outperforms existing methods, improving the compilation rate from 63% to 95.1% compared to the state-of-the-art SnR, marking a 50.1% improvement.
+On average, ZS4C can infer more accurate import statements (with an F1 score of 0.98) than SnR, with an improvement of 8.5% in the F1.
+Incomplete code snippets, Large Language Model, Program synthesis, API inference, Prompt engineering, ChatGPT
 1. Introduction
@@
 However, the majority of the code snippets on Q&A sites are uncompilable, which raises the obstacle for users to reuse or analyze these code snippets (Terragni et al., 2016; Rubei et al., 2020).
-Those code snippets are usually uncompilable and semantically incomplete for compilation due to unresolved types (e.g., what is the fully qualified name (FQN) for a type DateFormat) and missing dependent libraries (e.g., which library to import for DataFormat) (Terragni et al., 2016; Mishne et al., 2012; Dong et al., 2022).
+Those code snippets are usually uncompilable and semantically incomplete for compilation due to unresolved types (e.g., what is the fully qualified name (FQN) for the type DateFormat?) and missing dependent libraries (e.g., which library to import for DateFormat?) (Terragni et al., 2016; Mishne et al., 2012; Dong et al., 2022).
 Although missing such implementation details could be synthesized manually, it is tedious and often requires substantial familiarity with various libraries from developers.
 Furthermore, such manual synthesis effort is, however, not scalable for software engineering research based on Q&A sites’ code snippets (Zhang et al., 2021, 2019; Gao et al., 2015; Mishne et al., 2012).
+Besides, Integrated Development Environments (IDEs) sometimes fail to provide accurate import statements for the code snippets (JetBrains Community, 2018).
+This issue is also prevalent in LLM-generated code snippets, which can occasionally miss the appropriate import statements for successful execution (Pan et al., 2024).
 A technique to automatically repair code snippets by resolving the missing types and dependent libraries for compilation will significantly save developers time and facilitate the research based on such code snippets.
@@
 ZS4C operates in two stages.
-In the first stage, ZS4C utilizes LLM, i.e., ChatGPT, to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template.
+In the first stage, ZS4C utilizes LLM (e.g., ChatGPT), to identify missing import statements for a given code snippet, leveraging our designed task-specific prompt template.
 We apply the widely adopted self-consistency method to improve the robustness of the inference.
-In the second stage, ZS4C employs the Multi-Agent Conversation framework (Wu et al., 2023) to fix compilation errors caused by incorrect import statements and syntax errors.
-Specifically, ZS4C uses a compiler as an agent to provide feedback, i.e., compilation error logs, which serve as additional context for ChatGPT to correct the errors.
+In the second stage, ZS4C employs iterative conversations between an LLM and a validator (e.g., compiler) to fix errors caused by incorrect import statements and syntax errors.
+Specifically, ZS4C uses a validator to provide feedback, e.g., compilation error logs, which serve as additional context for an LLM to correct the errors.
 Compared with prior approaches, ZS4C is lightweight and does not require any training data, any pre-defined rules or heuristics, or program analysis like constructing AST.
-We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO consisting of 267 code snippets manually collected from Stack Overflow posts (Phan et al., 2018) against the SOTA approach SnR.
-ZS4C improves the compilation rate of SnR, from 63% to 87.6%, with a 39.3% improvement.
-On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 6.6% in the F1 score.
+We thoroughly evaluated ZS4C on a widely used benchmark called StatType-SO consisting of 267 code snippets manually collected from Stack Overflow posts (Phan et al., 2018) and a new dataset called Python-SO that contains 539 Python code snippets collected from Stack Overflow across 20 popular Python libraries.
+ZS4C outperforms baselines on all studied datasets.
+We evaluated ZS4C with two models, GPT-3.5 and GPT-4.
+ZS4C improves the compilation rate of the SOTA approach SnR, from 63% to 95.1%, with a 50.1% improvement.
+On average, ZS4C can infer more accurate import statements than SnR, with an improvement of 8.5% in the F1 score.
 Our ablation analysis shows that both components in ZS4C play a crucial role in the effectiveness of ZS4C.
-For instance, by employing the component conversation between a compiler and ChatGPT, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement.
+For instance, by employing the component conversation between a validator and GPT-3.5, ZS4C increases the compilation rate from 72% to 87.3% with a 21.3% improvement.
 Our parameter analysis reveals that ZS4C is not sensitive to parameters.
-Lastly, we conducted a failure analysis and our results sheds light on the limitations of using LLMs for synthesizing code, such as hallucination (e.g., fake FQN, and partially correct FQN) and unexpected code modification from LLMs.
+Lastly, we conducted a failure analysis and our results shed light on the limitations of using LLMs for synthesizing code, such as hallucination (e.g., fake FQN, and partially correct FQN) and unexpected code modification from LLMs.
 This paper makes the following contributions:
-  • We propose ZS4C, a novel zero-shot LLM-based approach to automatically synthesize compilable code for code snippets, which employs the multi-agent conversation framework to boost the successful rate of synthesizing compilable code for code snippets.
-  • We conducted extensive experiments to demonstrate the effectiveness of ZS4C compared with SOTA approach, and other baselines that mimic how developers use ChatGPT for resolving compilation issues for incomplete code snippets.
+  • We propose ZS4C, a novel zero-shot LLM-based approach to automatically synthesize compilable code for code snippets, which employs iterative conversations LLM and validator (e.g., compiler) to boost the successful rate of synthesizing compilable code for code snippets.
+  • We conducted extensive experiments to demonstrate the effectiveness of ZS4C compared with the SOTA approach, and other baselines that mimic how developers use LLMs (e.g., ChatGPT) for resolving compilation issues for incomplete code snippets.
   • We shed light on the limitation of using LLMs for the code synthesis task.
-  • We made our code and dataset publicly available (dat, 2023) to encourage future research on studying partial code synthesis and related approaches.
+  • We construct new data for Python libraries. We made our code and dataset publicly available (dat, 2023) to encourage future research on studying partial code synthesis and related approaches.
 2. Background and related work
@@
 In addition to these pre-trained code models, another research trajectory focuses on optimizing the application of LLMs to software engineering tasks such as code generation (Liu et al., [n.
-d.]), program repair (Wei et al., 2023), testing (Schäfer et al., 2023), and beyond.
+d.]), program repair (Wei et al., 2023; Liu et al., 2024a), testing (Schäfer et al., 2023), and beyond.
 Given the remarkable capability of LLMs to learn from zero or minimal examples, there’s a growing exploration into how in-context learning and prompt engineering can elevate LLM-based applications in software engineering.
@@
 They found that meticulously curated demonstrations can significantly outperform conventional demonstration construction techniques.
-Different from the aforementioned research, we focus on designing a new prompt engineering method to synthesize compilable code for incomplete code snippets on Q&A sites by leveraging LLM as a knowledge base.
+Liu et al.
+introduced the LLM-CompDroid framework, which combines the strengths of LLMs and traditional tools for repairing configuration compatibility bugs in Android apps (Liu et al., 2024a).
+Their experimental results demonstrate a significant enhancement in bug resolution performance and surpassing the state-of-the-art tool.
+Different from the aforementioned research, we focus on designing a new prompt engineering method to synthesize compilable code for incomplete code snippets on Q&A sites by leveraging LLM as a knowledge base and collaborating with a compiler.
 We introduce ZS4C, a lightweight zero-shot approach that decomposes the task into two stages, and enables a collaborative conversation between the LLM and a compiler to resolve compilation errors.
@@
 In Import Statement Inference, we leverage a large language model (LLM) to identify the missing dependent libraries (by identifying missing import statements) in the code snippet.
-In Conversational Error Fixing, we further fix compilation errors caused by the wrong dependent libraries inferred by Import Statement Inference and other compilation errors (e.g., syntax error) to boost the compilation success rate of a code snippet, through the conversation between a compiler and the LLM.
+In Conversational Error Fixing, we further fix compilation errors caused by the wrong dependent libraries inferred by Import Statement Inference and other compilation errors (e.g., syntax error) to boost the compilation success rate of a code snippet, through the conversation between a validator such as a compiler which could provide error log, and the LLM.
 For simplicity, in the following section, we refer to Import Statement Inference as ImportStInfer and Conversational Error Fixing as ConversationFixing.
@@
 We follow the style of prior work (Li et al., 2023; Xia and Zhang, 2023) in providing a short and concise prompt template tailored to the import statement inference task, as illustrated below.
-#Instruction Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Do not check for any import statements in the code.
+Prompt Template - 1 for Import Statement Inference #Instruction
+Presentation_Guide: “Reply with to-the-point answer, no elaboration.”
+Main_Instruction (for Java): “Do not check for any import statements in the code.
 Only give correct imports by not using wildcard imports.
 Please note that you need to pay close attention and your response should be specific and accurate.
-Avoid repetition and must not generate wrong and nonexistent imports:” #Query {$input\_code\_snippet$}
+Avoid repetition and must not generate wrong and nonexistent imports:”
+Main_Instruction (for Python): “Only give correct import statements for the attached code.
+Please note that you need to pay close attention and your response should be specific and accurate.
+Avoid repetition and must not generate wrong imports:”
+#Query
+{$input\_code\_snippet$}
 In Instruction, we first use Presentation_Guide “Reply with to-the-point answer, no elaboration” to guide the model to provide a compact and to-the-point answer to our query.
@@
 We design the instructions to guide and control the output of the model for our task step by step in Main_Instruction.
-We first ask the model not to check the code for any import statements if present.
+We first ask the model to ignore any import statements in the given input code if present.
 [^2] Note that, in line with previous studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), the input code snippets utilized for evaluation in this study do not contain any import statements.
+However, code snippets on SO sometimes come with import statements.
 This design directs the model’s attention exclusively to the body of the code rather than any existing import statements, which could be incorrect.
-Moreover, we observed that the model sometimes generates responses including “no imports are found”, likely stemming from a misinterpretation of the task.
+Moreover, we observed that the model sometimes generates responses including “no imports are found”, likely stemming from a misinterpretation of the task (e.g., to find import statements in the provided code snippet).
 To mitigate this issue, we explicitly instruct the model to avoid seeking import statements from the input code snippet.
-Next, we instruct the model to provide only correct imports and to avoid using wildcard imports.
-Wildcard import refers to the mechanism by which one can import all classes and interfaces from a particular package without specifying their fully qualified names by using the asterisk (*) character.
+Another issue is wildcard import, which refers to the mechanism by which one can import all classes and interfaces from a particular package without specifying their fully qualified names by using the asterisk (*) character.
 For instance, in Java, one can use “import java.util.*;” to import all classes from the “java.util” package, including the ArrayList, HashMap, and Scanner classes, among others, rather than individually importing classes, e.g., “import java.util.ArrayList”.
-We instruct the model to avoid wildcard imports for several reasons: 1) wildcard imports may cause naming conflicts if multiple classes with the same name exist in different packages; 2) we want to steer the model towards identifying the exact classes required rather than resorting to the short-cut wildcard imports; 3) wildcard imports are generally discouraged in practice because it can compromise the code readability and introduce unintended side-effects that can be very difficult to debug (Van Oort et al., 2021).
+Wildcard may cause several issues: 1) wildcard imports may cause naming conflicts if multiple classes with the same name exist in different packages; 2) we want to steer the model towards identifying the exact classes required rather than resorting to the short-cut wildcard imports; 3) wildcard imports are generally discouraged in practice because it can compromise the code readability and introduce unintended side-effects that can be very difficult to debug (Van Oort et al., 2021).
+Therefore, we instruct the model to provide only correct imports and to avoid using wildcard imports.
 Following the recommended best practice for prompt design mentioned in  (White et al., 2023), we add a sentence “Please note that you need to pay close attention and your response should be specific and accurate.” to explicitly emphasize the importance of specificity and accuracy for the LLM to prioritize.
-We also observed that in some cases, the model generates duplicate import statements and incorrect import statements that referenced non-existent libraries.
-To mitigate this issue, we added a sentence “Avoid repetition and must not generate wrong and nonexistent imports” to the prompt template.
+Without this sentence, LLM sometimes outputs explanations with the inferred import statements and brings us challenges in extracting the import statement from the output.
+We also observed that in some cases, the model generates duplicate import statements and incorrect import statements that reference non-existent libraries (e.g., hallucination).
+To mitigate this issue, we added the sentence “Avoid repetition and must not generate wrong and nonexistent imports” to the prompt template.
 Finally, we instantiate the {$input\_code\_snippet$} with the actual code snippet and send the filled prompt to LLM for inference.
@@
 If at least two lists have the same frequency, we repeat the process (i.e., running the same prompt for another $K$ times) until we break the tie.
-For instance, we query the model using the same prompt 5 times, and the model generates five lists, $L_{1}=\{A,B,C\}$, $L_{2}=\{B,A,C\}$, $L_{3}=\{A,B,C,D\}$, $L_{4}=\{A,B\}$, and $L_{5}=\{C,B,A\}$, we take $\{A,B,C\}$ as the final result as it (the combination of three libraries - A, B, C) appears 3 times with the highest frequency.
-In this study, we empirically set $K$ to 10, which achieves the best performance (see results of RQ4).
+For instance, if we query the model using the same prompt 5 times, and the model generates five lists, $L_{1}=\{A,B,C\}$, $L_{2}=\{B,A,C\}$, $L_{3}=\{A,B,C,D\}$, $L_{4}=\{A,B\}$, and $L_{5}=\{C,B,A\}$.
+We take $\{A,B,C\}$ as the final result as it (the combination of three libraries - A, B, C) appears 3 times with the highest frequency.
+In this study, we empirically set $K$ to 10, which achieves the best performance (see results of RQ3).
+Note that we only apply self-consistency in the Import Statement Inference component.
 3.2. Conversational Error Fixing
@@
 Drawing inspiration from recent advancements in the Multi-Agent Conversation and its applications across various tasks (Wu et al., 2023; Zhang et al., 2023; Xi et al., 2023), where multiple LLM-based agents work collaboratively and outperform a single LLM-based agent, we design a ConversationFixing component to further improve the results given by ImportStInfer.
-Within this component, a Java compiler is integrated into workflow with the LLM (via conversations) to address compilation errors.
-This compiler not only validates the generated code but also produces an error log that aids in resolving compilation issues.
-Specifically, we leverage the feedback, i.e., the error log, from the compiler’s validation as context, prompting the LLM to fix the errors for the given code snippet.
+Within this component, a validator (e.g., a compiler for Java code) is integrated into the workflow with the LLM (via conversations) to address compilation errors.
+This validator not only validates the generated code but also produces an error log that aids in resolving compilation issues.
+Specifically, we leverage the feedback, i.e., the error log, from the validator as context, prompting the LLM to fix the errors for the given code snippet.
+1
 Input : Code snippet with inferred import statements, $CS$
-Input : Maximum number of interaction between compiler and LLM, $M$
+Input : Maximum number of interaction between validator and LLM, $M$
 Input : Large language model, $LLM$
@@
 3 $history$ = ””
-3 // Conversation between compiler and LLM up to $K$ times
+4 // Conversation between validator and LLM up to $K$ times
 5 for $i\leftarrow 0$ to $M-1$ do
-5       // load dependent libraries dynamically
-6       Compiler.loadExternalLibraries($CS$)
-7       $CompileResults$ = Compiler.Compile($CS$)
-8       // $CS$ is failed to compile
-9       if $CompileResults$ is failed then
-10             $errorLog$ = getErrorLog(CompilerResults)
+6       $ValidationResults$ = Validator.validate($CS$)
+7       // $CS$ failed to pass the validator
+8       if $ValidationResults$ is failed then
+9             $errorLog$ = getErrorLog($ValidationResults$)
 10             $Prompt$ = constructPrompt($errorLog$, $history$)
 11             $CS$, $history$ = queryLLM($LLM$, $Prompt$)
-13
+12
 13      else
+14
 15            break
@@
 We present the logic of ConversationFixing in Algorithm 1.
-The algorithm takes the code snippet with inferred import statements generated by ImportStInfer, denoted as $CS$, the maximum number of interactions between compiler and LLM, represented by $M$, and a large language model, $LLM$, as input, and outputs a most recent fixed code, denoted as $FC$.
-First, we load external libraries and compile the $CS$ (lines 6-7) to examine if $CS$ can be compiled successfully.
-If the compilation is successful, we return $CS$ directly.
-If not, we construct a prompt incorporating the error log and $CS$ (lines 10 - 11).
-Subsequently, We query the $LLM$ with the constructed prompt, seeking corrections for $CS$ (line 12).
-This conversation (lines 5 - 15) is repeated up to $M$ times until a successful compilation.
-Note that during the compilation, the compiler needs to search for the external libraries to import as external dependencies (line 6).
+The algorithm takes the code snippet with inferred import statements generated by ImportStInfer, denoted as $CS$, the maximum number of interactions between validator and LLM, represented by $M$, and a large language model, $LLM$, as input, and outputs a most recent fixed code, denoted as $FC$.
+First, we examine if $CS$ can pass the validator successfully (line 5).
+If the validation is successful, we return $CS$ directly.
+If not, we construct a prompt incorporating the error log and $CS$ (lines 8 - 9).
+Subsequently, We query the $LLM$ with the constructed prompt, seeking corrections for $CS$ (line 10).
+This conversation (lines 4 - 14) is repeated up to $M$ times until a successful compilation.
+Prompt Template - 2 for Conversational Error Fixing Presentation_Guide: “Reply with to-the-point answer, no elaboration.”
+Main_Instruction (for Java): “Now fix the error by focusing on fixing the import statements by not using wildcard imports and must not modify code body which means do not change anything inside the class.
+So, it can be successfully compiled and reply with full code.”
+Main_Instruction (for Python): “Now fix the error by focusing on fixing the import statements. So, it can be run successfully and reply with full code.”
+# Prompt template for the 1st attempt:
+{Presentation_Guide}
+“See the code below:”
+$\{input\_code\_snippet\_with\_inferred\_import\_statements\}$
+“For the above code I got the below error log:”
+$\{error\_log\}$
+{Main_Instruction}
+# Prompt starting from the 2nd attempt:
+$Context$:
+$\{chat\_history\}$
+$\{last\_code\_from\_ChatGPT\}$
+“You gave the above imports fix in your attempt $\{attempt\_number\}$. But compiler gave this error:”
+$\{last\_error\_message\}$
+{Main_Instruction}
+In each iteration, we update the $CS$ as suggested by $LLM$ and provide the chat history as contextual information.
+If, after attempting $M$ times, $CS$ remains unsuccessful, the latest revised code is returned.
+Our design mimics how developers use an LLM (e.g., ChatGPT) to fix compilation errors for a piece of code over conversation with LLM (e.g., ChatGPT).
+We present our designed prompt template-2.
+We designed two templates: one for the first attempt and the other for subsequent attempt(s).
+In the first attempt, we attach the uncompilable code with an error log and ask LLM to fix the error and return a fixed version.
+Similar to the Prompt Template - 1 used in ImportStInfer (see Section 3.1), we ask LLM to avoid wildcard imports.
+We observed cases where the LLM circumvents errors by simply removing code segments that cause the compilation error.
+To mitigate this issue, we instruct the LLM to not modify the code body.
+For the following attempt, we attach the history from the previous attempts and ask LLM to fix based on history.
+Since the size of LLM’s context window is limited, we only keep the most recent chat history to offer the context that fits the window size.
+Note that during the validation, if a compiler is used as a validator for programming languages such as Java and C++ which needs compilation first, the compiler needs to search for the external libraries to import as external dependencies.
 To do so, we dynamically import the external libraries based on the import statements obtained from ImportStInfer or previous round of ConversationFixing.
 More specially, we construct a knowledge base to store the mappings between fully qualified names (FQN) of classes/functions and their corresponding external libraries (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar) in an offline manner.
-See more details of knowledge base construction in Section 4.2.
+See more details of knowledge base construction in Section 4.2.1.
 When a piece of code with inferred import statements is input into ConversationFixing, we search the external libraries for each FQN from the import statements and load the corresponding libraries from the constructed knowledge base.
@@
 We still compile the code and the error message generated by the compiler helps LLM fix the error (see more detailed results in Section 5.2).
-Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Now fix the error by focusing on fixing the import statements by not using wildcard imports and must not modify code body which means do not change anything inside the class.
-So, it can be successfully compiled and reply with full code.” # Prompt template for the 1st attempt: {Presentation_Guide} “See the code below:” $\{input\_code\_snippet\_with\_inferred\_import\_statements\}$ “For the above code I got the below error log:” $\{error\_log\}$Main_Instruction} # Prompt starting from the 2nd attempt: $Context$: $\{chat\_history\}$$\{last\_code\_from\_ChatGPT\}$ “You gave the above imports fix in your attempt $\{attempt\_number\}$.
-But compiler gave this error:” $\{last\_error\_message\}$Main_Instruction}
-In each iteration, we update the $CS$ as suggested by $LLM$ and provide the chat history as contextual information.
-If, after attempting $M$ times, $CS$ remains non-compilable, the latest revised code is returned.
-Our design mimics how developers use an LLM (e.g., ChatGPT) to fix compilation errors for a piece of code over conversation with LLM (e.g., ChatGPT).
-We present our designed prompt template-2.
-We designed two templates: one for the first attempt and the other for subsequent attempt(s).
-In the first attempt, we attach the uncompilable code with an error log and ask LLM to fix the error and return a fixed version.
-Similar to the Prompt Template - 1 used in ImportStInfer (see Section 3.1), we ask LLM to avoid wildcard imports.
-We observed cases where the LLM circumvents errors by simply removing code segments that cause the compilation error.
-To mitigate this issue, we instruct the LLM to not modify the code body.
-For the following attempt, we attach the history from the previous attempts and ask LLM to fix based on history.
-Since the size of LLM’s context window is limited, we only keep the most recent chat history to offer the context that fits the window size.
+For Java code snippets, we select the standard compiler Conversational Error Fixing component.
+For Python code snippets, we use Pyflakes (Developers, 2024), which is a static analysis tool for Python that checks for errors (syntax errors, undefined variables, and other common issues) in Python source files.
+For this study, we utilized Pyflakes version 3.2.0 to analyze our codebase before executing it with a Python interpreter.
 4. Experimental Design
@@
   • RQ2: What is the contribution of individual components to ZS4C?
-  • RQ3: Does few-shot learning improve ZS4C?
-  • RQ4: How do parameters affect the effectiveness of ZS4C?
-4.2. Datasets
-We use an established benchmark called StatType-SO, which was used in the related studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019).
+  • RQ3: How do parameters affect the effectiveness of ZS4C?
+  • RQ4: Does few-shot learning improve ZS4C?
+4.2. Data preparation
+4.2.1. Datasets
+To evaluate the effectiveness of ZS4C, we first use an established benchmark called StatType-SO, which was used in the related studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019).
 The dataset consists of 267 code snippets of six popular Java libraries.
-Table 1 summarizes the basic statistics of StatType-SO.
 The six libraries are of different sizes and cover various application domains.
@@
 This benchmark consists of the code snippets and their dependent libraries.
-For constructing the knowledge base that stores the mapping between fully qualified names (FQN) of classes/functions and their corresponding external libraries, we downloaded all versions of the most popular 500 libraries from maven repositories [^3] https://mvnrepository.com/popular.
-We extracted the FQN for all classes/functions from the jar files and created the knowledge base.
-Table 1. Basic statistics of StatType-SO.
-#Code Snippet #Average LOC #Unique Import Statements
-Android 50 38 129
-JDK 23 90 103
-Hibernate 5 50 123
-JodaTime 50 19 48
-GWT 50 35 108
-XStream 44 37 80
+To complement and diversify StatType-SO, which only consists of Java libraries, we also constructed a Python dataset following previous studies (Phan et al., 2018; Subramanian et al., 2014).
+We first selected the top 20 most popular Python libraries (Team, 2024).
+For each library $L$ of interest, we searched for all the posts from the past 10 years that use $L$ via Stack Overflow tags on StackExchange Data Explorer [^3] https://data.stackexchange.com/ and downloaded them.
+To ensure the quality of the posts, we only included those with a score greater than 5.
+We considered code snippets that have more than ten lines of code in both questions and answers.
+We ended up with 539 code snippets across the 20 studied libraries.
+To construct the benchmark (i.e., the correct import statements for each Python code snippet), the first author manually collected the required modules by examining the posts on Stack Overflow that contain the corresponding code snippet and searching on Google.
+Note that some code snippets initially came with partially imported modules; we removed those modules when constructing the dataset to avoid data leakage.
+Table 1 summarizes the basic statistics of StatType-SO and Python-SO.
+4.2.2. Knowledge base construction
+In Section 3.2, we use a compiler to validate the code snippet and provide feedback (i.e., error log) to the LLM for error fixing for Java code snippets.
+To make the compiler ready for compilation, we need to import proper external libraries for dependencies.
+For this, we need to construct the knowledge base that stores the mapping between fully qualified names (FQN) of classes/functions and their corresponding external libraries.
+To do so, we downloaded all versions of the most popular 500 libraries from maven repositories [^4] https://mvnrepository.com/popular.
+We then performed the following steps to extract the FQN and the last modification date for each external library.
+  • FQN Extraction: We first decompress the jar file for each external library.
+Next, we use the javap tool [^5] https://docs.oracle.com/javase/8/docs/technotes/tools/windows/javap.html, which is a Java class file disassembler, to analyze the binary class files extracted from the JAR.
+The output from javap is parsed to identify all non-abstract, non-interface class definitions, and then extract their fully qualified names.
+  • Lastest Modification Date: Note that we also need the modified date of each library to select the most recent library if multiple libraries meet constraints.
+To retrieve the last modification date of the JAR file, we examine the metadata associated with the file within the JAR.
+This date provides a timestamp indicating the most recent update or release of the JAR file.
+Note that for some jars we were not able to extract the valid modification date.
+Therefore, we added them manually from the Maven repository.
+We then create the inverse index for each pair of FQN and its corresponding library (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar).
+For Python code snippets, we do not use a compiler, since it is an interpreted language.
+Therefore, we did not construct a separate knowledge base.
+We use Pyflakes (Developers, 2024) to perform the validation and provide feedback in Section 3.2.
+After that, we execute Python code by invoking the Python interpreter as a subprocess within our scripts.
+To ensure a consistent and isolated environment for running the Python code, we use a virtual environment.
+We identified and listed the necessary libraries for Python-SO, ensuring their installation in the latest versions within this virtual environment.
+This strategy enabled us to manage dependencies efficiently and avoid conflicts with other projects.
+Table 1.
+Basic statistics of StatType-SO and Python-SO, including the number of code snippets, the average LOC, and unique import statement in each library.
+Note that #Unique Import Statements refers to the total number of unique import statements that need to be predicted correctly for each library.
+StatType-SO
+Library #Code Snippet Mean #LOC Max #LOC Standard Deviation of #LOC #Unique Import Statements
+Android 50 35 192 34 129
+JDK 23 82 300 64 103
+Hibernate 50 42 136 29 123
+JodaTime 50 17 49 7 48
+GWT 50 31 111 16 108
+XStream 44 35 100 18 80
+Python-SO
+Library #Code Snippet Mean #LOC Max #LOC Standard Deviation of #LOC #Unique Import Statements
+BeautifulSoup 35 13 29 6 31
+Flask 23 14 29 6 32
+Gensim 11 15 28 5 16
+Keras 17 24 116 24 38
+Lightgbm 7 18 36 11 13
+Matplotlib 49 13 36 7 18
+NLTK 18 14 46 9 26
+Numpy 17 22 109 24 21
+OpenCV 50 19 109 16 14
+Pandas 8 18 36 10 9
+Plotly 53 20 76 11 27
+Pytorch 26 16 30 7 24
+Scikit-learn 17 15 29 6 24
+Scipy 37 16 65 10 44
+Scrapy 49 19 113 18 74
+Seaborn 48 15 53 10 17
+Spacy 26 14 26 4 26
+Tensorflow 16 36 132 36 23
+Theano 11 11 18 3 20
+XGBoost 21 21 75 14 35
 4.3. Evaluation metrics
@@
 Criterion 1: Compilation Rate on the Synthesized Code.
-To measure the effectiveness of ZS4C in synthesizing compilable code from incomplete code snippets, we use compilation rate, i.e., the ratio of code snippets that successfully compile post-repair to the total number of snippets evaluated, following previous studies (Dong et al., 2022; Terragni et al., 2016).
+To measure the effectiveness of ZS4C in synthesizing compilable code from incomplete code snippets, we use compilation rate (CR), i.e., the ratio of code snippets that successfully compile post-repair to the total number of snippets evaluated, following previous studies (Dong et al., 2022; Terragni et al., 2016).
+Note that Python code does not need to be compiled, therefore, we use a standard Python Interpreter to execute the code, and if it passes, it typically refers to the successful execution.
+For simplicity, throughout the paper, we consider it compilable and refer to it as the compilation rate (CR).
 Criterion 2: Accuracy of the Synthesized Import Statements.
@@
 4.4. Base large language model
-In this study, we choose ChatGPT (OpenAI, 2023a) as our base LLM.
-We select ChatGPT as it is recognized to be one of the most powerful LLMs and has gained wide attention in recent months by demonstrating the outstanding capability of solving tasks in various areas (Frieder et al., 2023), including software engineering (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023; Xia and Zhang, 2023).
-We accessed ChatGPT, i.e., the default GPT-3.5-Turbo model, via the OpenAI API (OpenAI, 2023b).
+In this study, we choose GPT-3.5 and GPT-4 as our base LLMs (OpenAI, 2023a).
+We select them as they are recognized to be the most powerful LLMs and have gained wide attention in recent months by demonstrating the outstanding capability of solving tasks in various areas (Frieder et al., 2023), including software engineering (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023; Xia and Zhang, 2023).
+We accessed the “gpt-3.5-turbo-0125” for GPT-3.5 and “gpt-4o” model for GPT-4, via the OpenAI API (OpenAI, 2023b, c).
+We denote the ZS4C with GPT-3.5 and GPT-4 as ZS4CGPT-3.5 and ZS4CGPT-4, respectively.
 4.5. Implementation details
@@
 We leveraged the replication package of SnR as provided in their paper.
-[^4] We found disparity between the number reported in the paper and the number derived from their replication package.
+[^6] We found the disparity between the number reported in the paper and the number derived from their replication package.
 We present the numbers derived from their replication package in this paper.
 We do not compare with CSNIPPEx because this Eclipse plugin is no longer accessible.
-  • BasePrompt: We also consider a baseline that employs a straightforward prompt to instruct a LLM to make the given code snippet compilable.
+  • BasePrompt: We also consider a baseline that employs a straightforward prompt to instruct an LLM to make the given code snippet compilable.
 Specifically, we replace the Main_Instruction as “Make the code below compilable: {$input\_code\_snippet$}” in Prompt Template-1 (ref.
 Section 3.1) and while retaining the remainder of the prompt template unchanged.
-We evaluate the effectiveness of ZS4C against the two baselines on the two criteria outlined in Section 4.3.
+We denote the BasePrompt with GPT-3.5 and GPT-4 as BasePromptGPT-3.5 and BasePromptGPT-4, respectively.
+We evaluate the effectiveness of ZS4C with GPT-3.5 and GPT4 against the two baselines on the two criteria outlined in Section 4.3 on StatType-SO.
+Since SnR is designed for Java and is not appliable to Python code, we focus on comparing ZS4C with BasePrompt for Python-SO dataset.
 To ensure the reliability of our results and mitigate potential biases arising from the inherent randomness of the LLM, we run both ZS4C and BasePrompt five times and take the median value as the final result.
 4.6.2. Approach of RQ2
-[…]
+In this RQ, we carry out an ablation study to understand the contribution of each component (i.e., ConversationFixing and ImportStInfer) in ZS4C.
+Note that for ConversationFixing, the input code snippet will be the original code snippet without any inferred import statements.
+We are also interested in understanding the contribution of the self-consistency method using in ImportStInfer.
+Therefore, we remove self-consistency from ImportStInfer (refer as to ImportStInfernoSC) and compare it with ImportStInfer.
+We then evaluate the effectiveness of the above three approaches following the same methodology in RQ1, and compare them with ZS4C.
+Note that using the API of GPT-4 is expensive, and we observe similar performance between GPT-4 and GPT-3.5.
+Therefore, we carry out the ablation study with the more economical model GPT-3.5.
 4.6.3. Approach of RQ3
+ZS4C has two parameters: $K$ the number of sampling in self-consistency, and $M$ the maximum rounds of conversation between the compiler and the LLM in Conversational Error Fixing.
+In this RQ, we investigate the impact of each parameter on the effectiveness of ZS4C, where we set $K$ to 10, 20, and 30 and set $M$ to 5, 10, and 15, respectively.
+Similar to RQ2, we carry out the ablation study with GPT-3.5.
+4.6.4. Approach of RQ4
 Few-shot in-context learning is a paradigm where a limited set of examples are provided in the prompt and allows an LLM to learn from the examples in a specific context.
@@
 Based on the code representations output by CodeBERT, we select the code snippets presenting the highest cosine similarities with the input code snippet.
-Due to the window size token limitation of ChatGPT (i.e., 4k), we only put the most similar code snippet and its import statements in the prompt as the demonstration example.
+When testing the performance of ZS4C with few-shot in-context learning, we selected different configurations based on the capabilities of the language models and their token window size limitations.
+Table 2 presents the five-number summary of the prompt token size for different settings.
+As we can see the median size for three-shot prompts is 7,100, which indicates that the majority of the prompts exceed the window size of GPT-3.5.
+Therefore, for GPT-3.5, we utilized a one-shot learning approach, while for GPT-4, we employed both one-shot and three-shot learning approaches.
+This decision was made due to the token window size constraints (i.e., 4k input tokens for GPT-3.5 and 128k input tokens for GPT-4).
 We put the retrieved demonstrated example after the Main_Instruction in Prompt Template - 1 as shown in Section 3.1.1.
-In other words, we create a one-shot version of ZS4C.
-4.6.4. Approach of RQ4
-ZS4C has two parameters: $K$ the number of sampling in self-consistency, and $M$ the maximum rounds of conversation between the compiler and the LLM in Conversational Error Fixing.
-In this RQ, we investigate the impact of each parameter on the effectiveness of ZS4C, where we set $K$ to 10, 20, and 30 and set $M$ to 5, 10, and 15, respectively.
+Table 2.
+Summary of prompt size for different settings in terms of input token size distributions, including the five-number summary statistics (Min: Minimum, Q1: 25th percentile, Median: 50th percentile, Q3: 75th percentile, Max: Maximum).
+StatType-SO Python-SO
+Min Q1 Median Q3 Max Min Q1 Median Q3 Max
+Three-shot 4,920 5,654 7,100 10,380 100,650 2,370 4,200 5,790 8,430 42,060
+One-shot 1,200 1,934 3,320 4,840 65,020 1,580 2,800 3,860 5,620 28,040
+Zero-shot 620 1,082 1658 2,440 32,530 790 1,400 1,930 2,810 14,020
 5. Results
-5.1. Results of RQ1
+5.1. RQ1: How effective is our approach ZS4C?
 Table 3.
-The comparison of SnR, BasePrompt, and ZS4C in terms of F1, precision (Pre), and recall (Rec), and compilation rate (CR).
+The comparison of SnR, BasePrompt, ZS4C with GPT-3.5 and ZS4C with GPT-4 for StatType-SO (Java) dataset in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR).
 We present the average for F1, Rec, Pre, and the total compilation rate in the summary row.
 The cells that have the best performance are marked in bold.
-SnR BasePrompt ZS4C
-F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
-Android
-0.952
-0.909
-1
-37
-0.967
-0.95
-0.987
-44
-0.990
-0.990
-0.990
-48
-JDK
-0.887
-0.796
-1
-12
-0.482
-0.317
-1
-15
-0.997
-1.000
-0.994
-23
-Hibernate
-0.948
-0.918
-0.980
-23
-0.75
-0.606
-0.985
-22
-0.902
-0.834
-0.982
-32
-Joda-Time
-0.860
-0.759
-0.992
-37
-0.979
-0.971
-0.988
-44
-0.985
-0.994
-0.977
-50
-GWT
-0.835
-0.721
-0.991
-30
-0.951
-0.927
-0.977
-32
-0.968
-0.940
-0.997
-46
-XStream
-0.934
-0.877
-1
-29
-0.863
-0.78
-0.967
-22
-0.934
-0.890
-0.980
-35
-Summary
-0.903
-0.830
-0.994
-168 (62.9%)
-0.832
-0.759
-0.984
-179 (67%)
-0.963 0.938
-0.987
-234 (87.6%)
-[…]
-Same Different Extra Missing None
-SnR 143 78 4 35 7
-BasePrompt 173 55 14 19 6
-ZS4C 202 35 16 14 0
-ImportStInfer 191 48 14 12 2
-[…]
-ConversationFixing 175 48 22 19 3
-We present the comparison results among ZS4C and two baseline approaches in Tables 2 and  3.
-The experimental results show that ZS4C successfully synthesized 234 compilable code snippets out of the 267 provided incomplete code snippets, resulting in an impressive overall compilation rate of 87.6%, outperforming two baselines.
-Compared with SnR, ZS4C showed a 39.3% improvement in terms of compilation rate.
+SnR BasePromptGPT-3.5 BasePromptGPT-4 ZS4CGPT-3.5 ZS4CGPT-4
+F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
+Android 0.952 0.909 1 37 (74%) 0.967 0.95 0.987 44 (88%) 0.969 0.956 0.983 45 (90%) 0.990 0.990 0.990 48 (96%) 0.988 0.99 0.987 48 (96%)
+JDK 0.887 0.796 1 12 (52.2%) 0.482 0.317 1 15 (65.2%) 0.713 0.557 0.989 18 (78.3%) 0.997 1.000 0.994 23 (100%) 0.994 1 0.988 23 (100%)
+Hibernate 0.948 0.918 0.98 23 (46%) 0.75 0.606 0.985 22 (44%) 0.834 0.747 0.944 29 (58%) 0.902 0.834 0.982 32 (64%) 0.956 0.916 1 41 (82%)
+Joda-Time 0.86 0.759 0.992 37 (74%) 0.979 0.971 0.988 44 (88%) 0.991 0.994 0.988 49 (98%) 0.985 0.994 0.977 50 (100%) 0.994 0.988 1 50 (100%)
+GWT 0.835 0.721 0.991 30 (60%) 0.951 0.927 0.977 32 (64%) 0.906 0.889 0.924 42 (84%) 0.968 0.94 0.997 46 (92%) 0.987 0.978 0.997 50 (100%)
+XStream 0.934 0.877 1 29 (65.9%) 0.863 0.78 0.967 22 (50%) 0.927 0.89 0.967 38 (86.4%) 0.934 0.89 0.98 35 (79.6%) 0.959 0.925 0.995 42 (95.5%)
+Summary 0.903 0.83 0.994 168 (62.9%) 0.832 0.759 0.984 179 (67%) 0.890 0.839 0.966 221 (82.8%) 0.963 0.938 0.987 234 (87.6%) 0.98 0.966 0.995 254 (95.1%)
+Table 4.
+The comparison of BasePrompt (with GPT-3.5 and GPT-4) and ZS4C (with GPT-3.5 and GPT-4) for Python dataset (Python-SO) in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR).
+We present the average for F1, Rec, Pre, and the total compilation rate in the summary row.
+The cells that have the best performance are marked in bold.
+BasePromptGPT-3.5 BasePromptGPT-4 ZS4CGPT-3.5 ZS4CGPT-4
+F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
+BeautifulSoup 0.954 0.913 1 29 (82.9%) 0.947 0.9 1 31 (88.6%) 0.968 0.938 1 31 (88.6%) 0.974 0.95 1 33 (94.3%)
+Flask 0.937 0.908 0.967 15 (65.2%) 0.977 0.985 0.97 18 (78.3%) 1 1 1 23 (100%) 1 1 1 23 (100%)
+Gensim 0.919 0.895 0.944 10 (90.9%) 0.842 0.842 0.842 9 (81.8%) 1 1 1 11 (100%) 0.973 0.947 1 11 (100%)
+Keras 0.947 0.953 0.942 15 (88.2%) 0.948 0.965 0.932 13 (76.5%) 0.988 0.988 0.988 14 (82.4%) 0.994 0.988 1 15 (88.2%)
+LightGBM 0.947 0.9 1 6 (85.7%) 0.976 1 0.952 6 (85.7%) 1 1 1 7 (100%) 1 1 1 7 (100%)
+Matplotlib 0.983 0.978 0.989 49 (100%) 0.989 0.989 0.989 49 (100%) 0.994 0.989 1 46 (93.9%) 0.994 0.989 1 47 (95.9%)
+NLTK 0.968 0.957 0.978 10 (55.6%) 0.957 0.957 0.957 13 (72.2%) 1 1 1 18 (100%) 0.978 0.957 1 18 (100%)
+Numpy 0.887 0.796 1 15 (88.2%) 0.887 0.796 1 17 (100%) 0.826 0.704 1 16 (94.1%) 0.839 0.722 1 17 (100%)
+OpenCV 0.984 0.989 0.978 25 (50%) 0.989 1 0.979 31 (62%) 1 1 1 48 (96%) 1 1 1 49 (98%)
+Pandas 0.973 0.947 1 7 (87.5%) 0.947 0.947 0.947 8 (100%) 0.973 0.947 1 7 (87.5%) 0.973 0.947 1 7 (87.5%)
+Plotly 0.92 0.886 0.956 24 (45.3%) 0.946 0.919 0.974 27 (50.9%) 0.94 0.887 1 40 (75.5%) 0.937 0.896 0.982 42 (79.2%)
+Pytorch 0.945 0.968 0.923 23 (88.5%) 0.903 0.903 0.903 25 (96.2%) 0.984 0.968 1 26 (100%) 0.949 0.903 1 26 (100%)
+Scikit-learn 1 1 1 17 (100%) 1 1 1 17 (100%) 1 1 1 17 (100%) 0.989 0.979 1 17 (100%)
+Scipy 0.963 0.948 0.979 24 (64.9%) 0.958 0.938 0.978 26 (70.3%) 0.974 0.949 1 37 (100%) 0.979 0.959 1 37 (100%)
+Scrapy 0.875 0.875 0.875 32 (65.3%) 0.877 0.892 0.863 34 (69.4%) 0.966 0.934 1 47 (95.9%) 0.97 0.942 1 49 (100%)
+Seaborn 0.974 0.974 0.974 45 (93.7%) 0.965 0.974 0.957 47 (97.9%) 0.987 0.974 1 48 (100%) 0.987 0.974 1 48 (100%)
+Spacy 0.926 0.875 0.982 18 (69.2%) 0.935 0.906 0.967 25 (96.2%) 0.984 0.969 1 16 (61.5%) 0.984 0.969 1 23 (88.5%)
+Tensorflow 0.961 0.961 0.961 7 (43.8%) 0.949 0.922 0.979 10 (62.5%) 0.99 0.98 1 13 (81.3%) 1 1 1 15 (93.8%)
+Theano 0.972 0.946 1 9 (81.8%) 0.946 0.946 0.946 8 (72.7%) 0.974 0.974 1 8 (72.7%) 0.986 0.973 1 10 (90.9%)
+XGBboost 0.971 0.944 1 17 (81%) 0.965 0.958 0.971 18 (85.7%) 0.986 0.972 1 18 (85.7%) 0.986 0.972 1 20 (95.2%)
+Summary 0.95 0.931 0.972 397 (73.7%) 0.945 0.937 0.955 432 (80.2%) 0.977 0.959 0.999 491 (91.1%) 0.975 0.953 0.999 514 (95.4%)
+Table 5. The comparison of SnR, BasePrompt, ZS4C in terms of list-wise matching degree on StatType-SO and Python-SO datasets.
+StatType-SO
+Approach Same Different Extra Missing None
+SnR 143 (53.56%) 78 (29.21%) 4 (1.5%) 35 (13.11%) 7 (2.62%)
+BasePromptGPT-3.5 173 (64.79%) 55 (20.60%) 14 (5.24%) 19 (7.12%) 6 (2.25%)
+BasePromptGPT-4 226 (84.64%) 24 (8.99%) 10 (3.75%) 7 (2.62%) 0 (0%)
+ZS4CGPT-3.5 202 (75.66%) 35 (13.11%) 16 (5.99%) 14 (5.24%) 0 (0%)
+ZS4CGPT-4 226 (84.64%) 24 (8.99%) 10 (3.75%) 7 (2.62%) 0 (0%)
+Python-SO
+Approach Same Different Extra Missing None
+BasePromptGPT-3.5 447 (82.93%) 43 (7.98%) 27 (5.01%) 19 (3.53%) 3 (0.56%)
+BasePromptGPT-4 447 (82.93%) 42 (7.79%) 39 (7.24%) 9 (1.67%) 2 (0.37%)
+ZS4CGPT-3.5 491 (91.09%) 35 (6.49%) 4 (0.74%) 9 (1.67%) 0 (0%)
+ZS4CGPT-4 499 (92.58%) 30 (5.57%) 3 (0.56%) 7 (1.30%) 0 (0%)
+We present the comparison results for StatType-SO in Tables 3 and for Python-SO in Table 4.
+The results show that both ZS4CGPT-3.5 and ZS4CGPT-4 outperform the baseline SnR by a large margin on StatType-SO.
+For instance, ZS4CGPT-3.5 successfully synthesized 234 compilable code snippets out of the 267 provided incomplete code snippets, resulting in an impressive overall compilation rate of 87.6%, outperforming two baselines.
+Compared with SnR, ZS4CGPT-3.5 showed a 39.3% improvement in terms of compilation rate.
 [^7] SnR paper reports that they can compile 197 out of 267 (73.8%) code snippets successfully after the repair.
@@
 Nevertheless, our approach outperforms SnR by improving the number of compilable code snippets to 234.
-Furthermore, ZS4C consistently outperformed the two approaches for all libraries.
-A major factor behind this boosting in compilation rate is the ConversationFixing component, which enables conversations between the compiler and the LLM.
-A deeper dive into the effectiveness and impact of this component can be found in RQ2 and Section 6.1.
-When measuring the accuracy of the synthesized import statements, on average, ZS4C achieves 0.961, 0.987, and 0.938 in terms of F1, Precision, and Recall, respectively, as detailed in Table 2.
-Compared with SnR, ZS4C achieves an improvement of 6.6% and 13.4% in terms of F1 and precision.
-In terms of Recall, ZS4C has a slightly lower performance than SnR, which achieves an almost perfect precision of 0.994 due to its constraint-based inference.
-In six of the target libraries, ZS4C outperforms SnR in five, except Hibernate.
+Furthermore, ZS4CGPT-3.5 consistently outperformed the two approaches for all libraries.
+ZS4CGPT-4 achieves a compilation performance improvement of 8.56% over its predecessor GPT-3.5, increasing from 87.6% to 95.1%.
+Similarly, on Python-SO, ZS4C outperforms BasePrompt.
+For instance, when using GPT3.5 as the base model, ZS4C achieves a compilation rate of 91.1%
+When measuring the accuracy of the synthesized import statements, ZS4C always outperforms baselines in all evaluation metrics.
+For instance, on average, ZS4CGPT-4 achieves 0.98, 0.966, and 0.995 in terms of F1, Precision, and Recall on StatType-SO, respectively, as detailed in Table 3.
+Compared with SnR, ZS4CGPT-4 achieves an improvement of 8.5% and 16.4% in terms of F1 and recall.
+ZS4CGPT-4 and SnR have similar precision.
+In six of the target libraries, ZS4CGPT-4 outperforms SnR in all six libraries.
 We perform failure cases analysis to understand why eventually some cases are not compilable after the fixing (see more details in Section 6.2).
-Compared with BasePrompt, ZS4C can compile 55 more code snippets, reflecting a 30.7% enhancement in the compilation rate.
-Furthermore, ZS4C elevates the accuracy of import statement inference, with the F1 score rising from 0.832 to 0.961.
-These improvements indicate the significance of task-specific prompt engineering in optimizing the performance of LLMs like ChatGPT.
-To complement accuracy metrics, we also compare the list-wise matching degree.
-ZS4C (202) can infer more exact match import statement lists (same) than SnR (143) [^6] SnR paper reports that they can achieve 183 same, 62 different, 11 extra, 4 missing, and 7 none matches.
+Compared with BasePrompt, ZS4C consistently can predict import statements more accurately and compile more code snippets.
+For instance, on StatType-SO ZS4CGPT-3.5 can compile 55 more code snippets, reflecting a 30.7% enhancement in the compilation rate, compared with BasePromptGPT-3.5.
+Furthermore, ZS4CGPT-3.5 elevates the accuracy of import statement inference, with the F1 score rising from 0.832 to 0.961.
+These improvements indicate the significance of task-specific prompt engineering in optimizing the performance of LLMs.
+We observe a similar trend on Python-SO dataset.
+ZS4C consistently outperforms BasePrompt with the same base model.
+To complement accuracy metrics, we also compare the list-wise matching degree and present the result in Table 5.
+Both ZS4CGPT-4 (226) and ZS4CGPT-3.5 (202) can infer more exact match import statement lists (same) than SnR  [^8] SnR paper reports that they can achieve 183 same, 62 different, 11 extra, 4 missing, and 7 none matches.
 However, in their replication package, we found the number we reported in the table.
-Nevertheless ZS4C can achieve more exact same matches.
-and BasePrompt (193).
-Also ZS4C can always output thing that is related to the correct import statements (none is 0).
+Nevertheless ZS4C with GPT-3.5 can achieve more exact same matches.
+and BasePrompt (193) on StatType-SO.
+We observe a similar trend on Python-SO, ZS4C always outperforms BasePrompt.
+Also ZS4C can always output things that are related to the correct import statements (none is 0).
 To understand why ZS4C outperforms SnR, we manually study the cases where ZS4C performs better than SnR.
+Since ZS4CGPT-4 and ZS4CGPT-3.5 achieve similar performance, here we study the cases from ZS4CGPT-3.5.
 We observe two reasons.
@@
 Therefore, in certain cases, SnR failed in the first repairing step.
-While ZS4C does not rely on static analysis (e.g., parsing AST).
+While ZS4C with GPT-3.5 does not rely on static analysis (e.g., parsing AST).
 LLM directly takes code snippets as sequence input and does the inference, which reduces the risk of failures due to errors incurring in static analysis.
@@
 While ZS4C leverages LLMs as the knowledge base to perform the inference, LLMs have sufficient knowledge to decide which libraries are most appreciated given the context.
-ZS4C improves the compilation rate of the SOTA approach, SnR, from 63% to 87.6%.
-On average, ZS4C’s synthesized import statements are more accurate than SnR, with an improvement of 6.6% in the F1 score.
-Furthermore, ZS4C is more effective than BasePrompt, which represents a straightforward way of utilizing ChatGPT for this task.
+In addition, we observe that ZS4CGPT-4 consistently outperforms ZS4CGPT-3.5 or achieves comparable effectiveness as ZS4CGPT-3.5 on both studied datasets, typically in terms of compilation rate.
+ZS4CGPT-4 improves the compilation rate of ZS4CGPT-3.5 from 87.6% to 95.1%.
+This probably originates from two advantages of GPT-4.
+First, GPT-4 is more advanced than GPT-3.5 and has already proved more effective in various code-related tasks (Liu et al., 2024b; Poldrack et al., 2023).
+Second, GPT-4 has a long window size and can fit much more context in the prompt (e.g., history generated during Conversational Error Fixing).
+In general, ZS4C with GPT-4 outperforms GPT-3.5 and other baselines on both StatType-SO and Python-SO.
+For instance, on StatType-SO, ZS4C with GPT-4 improves the compilation rate of the SOTA approach, SnR, from 63% to 95.1%.
+On average, ZS4C’s (with GPT-4) synthesized import statements are more accurate (with an F1 score of 0.98) than SnR, with an improvement of 8.5% in the F1 score.
+Furthermore, ZS4C is more effective than BasePrompt.
 5.2. RQ2: What is the contribution of individual components to ZS4C?
-[…]
-[…]
+Table 6.
+The comparison of ImportStInfernoSC, ImportStInfer, ConversationFixing in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR) with GPT-3.5.
+We present the average for F1, Rec, and Pre and the total compilation rate in the summary row.
+ImportStInfernoSC ImportStInfer ConversationFixing
 F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
-Android
-0.948
-0.915
-0.982
-39
-0.980
-0.980
-0.980
-42
-0.927
-0.902
-0.954
-43
-JDK
-0.919
-0.850
-1
-19
-0.997
-1
-0.994
-23
-0.825
-0.707
-0.992
-21
-Hibernate
-0.886
-0.813
-0.974
-27
-0.898
-0.828
-0.981
-38
-0.876
-0.803
-0.963
-34
-Joda-Time
-0.964
-0.941
-0.988
-42
-0.967
-0.959
-0.976
-29
-0.968
-0.965
-0.970
-49
-GWT
-0.939
-0.905
-0.976
-35
-0.956
-0.924
-0.990
-45
-0.888
-0.879
-0.896
-44
-XStream
-0.884
-0.936
-0.837
-31
-0.910
-0.850
-0.980
-28
-0.898
-0.850
-0.951
-34
-Summary
-0.923
-0.893
-0.960
-193 (72.2%)
-0.951
-0.924
-0.984
-205 (76.8%)
-0.897
-0.851
-0.954
-225 (84.2%)
-[…]
-We additionally observe that ConversationFixing achieves a higher compilation rate (84.2%) than ImportStInfer (72.2%).
+StatType-SO 0.923 0.893 0.960 193 (72.2%) 0.951 0.924 0.984 205 (76.8%) 0.897 0.851 0.954 225 (84.2%)
+Python-SO 0.956 0.937 0.978 417 (77.4%) 0.965 0.942 0.993 451 (83.7%) 0.942 0.908 0.983 463 (85.9%)
+Table 6 presents the results of the ablation analysis of ZS4C with GPT-3.5 on StatType-SO and Python-SO datasets.
+The experimental results show that the self-consistency method can improve the effectiveness of the Import Statement Inference component.
+Specifically, ImportStInfer has a 6.4% and 2% higher of compilation rate than ImportStInfernoSC on StatType-SO and Python-SO, respectively.
+The self-consistency method improves the results of ImportStInfernoSC from 0.923 to 0.951 and 0.956 to 0.965 in terms of average F1, on StatType-SO and Python-SO, respectively.
+We additionally observe that ConversationFixing achieves a higher compilation rate than ImportStInfer on both datasets.
 This underscores the benefits derived from the collaborative conversations between the LLM and the compiler.
@@
 Yet, when we shift the focus to the efficacy of both components in inferring missing import statements, the tables turn.
-ImportStInfer has a F1 of 0.951, which is much higher than ConversationFixing’s F1 score of 0.897.
+On both datasets, ImportStInfer has a higher F1 score than ConversationFixing.
 This demonstrates that extracting all required important statements from the code body can be difficult when the primary goal of the LLM is to resolve compilation errors.
@@
 Additionally, the self-consistency method improves the effectiveness of the Import Statement Inference component.
-5.3. RQ3: Does few-shot learning improve ZS4C?
-Table 5 presents the results of ZS4C in two settings, zero-shot and one-shot.
-Surprisingly, one-shot in-context learning does not outperform zero-shot in any evaluation metrics.
-Instead, there is a slight decline in terms of F1.
-We suspect the reason could have originated from the advantages of using demonstration examples.
-Demonstration examples are usually provided for in-context learning for two purposes: 1) to help LLM understand the intent of a task and learn the patterns to complete the task (White et al., 2023; Webson and Pavlick, 2021; Lu et al., 2021); 2) to provide domain-specific knowledge for downstream tasks (Nashid et al., 2023; Gao et al., 2023; Zhou et al., 2022a).
-In our cases, apparently, none of the above advantages of providing examples hold.
-Our task is straightforward, and the model can understand the intent expressed in the zero-shot prompt precisely.
-This is possible because ChatGPT is trained on a large dataset Common Crawl, which includes data from Stack Overflow and Github (Nagel, 2022).
-ChatGPT could understand the task precisely.
-Meanwhile, the demonstrations do not reveal any domain-specific knowledge that can be utilized.
-Table 5. The comparison of the performance of ZS4C in one-shot and zero-shot settings.
-Compilation Rate F1 Recall Precision
-one-shot 86.9% 0.955 0.934 0.981
-zero-shot 87.6% 0.961 0.938 0.987
-Surprisingly, one-shot in-context learning does not improve the performance of ZS4C.
-5.4. RQ4: How do parameters affect the effectiveness of ZS4C?
-Figure 3. The impact of parameters $K$ (left) and $M$ (right) on ZS4C.
+5.3. RQ3: How do parameters affect the effectiveness of ZS4C?
+Figure 3. The impact of parameters $K$ (left) and $M$ (right) on ZS4CGPT-3.5 for StatType-SO.
 Figure 3 (left) presents the results of parameter sensitivity analysis on $K$ (i.e., number of sampling in self-consistency).
-As $K$ increases, the effectiveness of ZS4C to infer the import statements declines slightly, from 0.961 to 0.934 in terms of F1.
+As $K$ increases, the effectiveness of ZS4C with GPT-3.5 for StatType-SO (Java) dataset to infer the import statements declines slightly, from 0.961 to 0.934 in terms of F1.
 We can observe similar trends for accuracy, recall, and precision.
@@
 Overall, the effectiveness of ZS4C is not sensitive to its parameters.
+5.4. RQ4: Does few-shot learning improve ZS4C?
+Table 8 presents the results of ZS4C in three settings, zero-shot, one-shot, and three-shot (few-shot) of ZS4CGPT-3.5 and ZS4CGPT-4 on both datasets.
+Surprisingly, one-shot in-context learning does not outperform zero-shot with GPT-3.5 in any evaluation metrics.
+Instead, there is a slight decline in terms of F1.
+With GPT-4, we observe a similar trend, where few-shot in-context learning does not improve the performance of ZS4C.
+Instead, the performance declines in the three-shot setting compared with zero-shot in terms of CR on StatType-SO.
+One possibility is that the code retrieved is not similar to the input code so the few-shot setting does not help.
+To verify this, we examine the similarity between each input code and their corresponding retrieved code snippet.
+Table 7 presents the similarity score and we observe that the retrieved code is actually similar to the input code snippet.
+So this assumption is rejected.
+Then, we suspect the reason could be that the demonstration examples do not provide any additional information to the model.
+Demonstration examples are usually provided for in-context learning for two purposes: 1) to help LLM understand the intent of a task and learn the patterns to complete the task (White et al., 2023; Webson and Pavlick, 2021; Lu et al., 2021); 2) to provide domain-specific knowledge for downstream tasks (Nashid et al., 2023; Gao et al., 2023; Zhou et al., 2022a).
+In our cases, apparently, none of the above advantages of providing examples hold.
+The model can understand the intent expressed in the zero-shot prompt accurately because of the clear problem definition.
+The reason may be that GPT is large enough to understand many programming tasks precisely.
+We demonstrate that with a powerful model, zero-shot synthesis of compilable code for incomplete code snippets using LLMs is feasible.
+Table 7. Summary of similarity score between each input code and their corresponding retrieved code snippet for one-shot setting in StatType-SO and Python-SO datasets.
+StatType-SO
+Library Minimum Mean Maximum
+Android 0.901 0.962 0.985
+JDK 0.791 0.924 0.973
+Hibernate 0.912 0.95 0.984
+JodaTime 0.937 0.963 0.986
+GWT 0.935 0.965 0.984
+XStream 0.938 0.953 0.983
+Python-SO
+Library Minimum Mean Maximum
+BeautifulSoup 0.912 0.953 0.98
+Flask 0.914 0.952 0.984
+Gensim 0.904 0.931 0.945
+Keras 0.867 0.945 0.985
+Lightgbm 0.922 0.944 0.958
+Matplotlib 0.816 0.932 0.989
+NLTK 0.876 0.941 0.972
+Numpy 0.85 0.932 0.968
+OpenCV 0.81 0.954 0.986
+Pandas 0.887 0.951 0.988
+Plotly 0.928 0.964 0.986
+Pytorch 0.912 0.937 0.956
+Scikit-learn 0.894 0.94 0.986
+Scipy 0.845 0.937 0.981
+Scrapy 0.869 0.957 0.983
+Seaborn 0.727 0.949 0.988
+Spacy 0.695 0.936 0.97
+Tensorflow 0.848 0.949 0.985
+Theano 0.891 0.938 0.951
+XGBoost 0.916 0.958 0.988
+Table 8. The comparison of the performance of ZS4C in one-shot and zero-shot settings. Here, CR means compilation rate.
+StatType-SO Python-SO
+CR F1 Recall Precision CR F1 Recall Precision
+Zero-shot (GPT-3.5) 87.6% 0.961 0.938 0.987 91.1% 0.977 0.959 0.999
+One-shot (GPT-3.5) 86.9% 0.955 0.934 0.981 90.2% 0.958 0.942 0.979
+Zero-shot (GPT-4) 95.1% 0.98 0.966 0.995 95.2% 0.975 0.953 0.999
+One-shot (GPT-4) 95.9% 0.985 0.974 0.999 94.8% 0.976 0.96 0.995
+Three-shot (GPT-4) 93.6% 0.986 0.974 0.999 95.4% 0.973 0.963 0.996
+Surprisingly, for both models GPT-3.5 and GPT-4, few-shot in-context learning does not improve the performance of ZS4C.
+Typically, with GPT-4, ZS4C can achieve a high F1 and Compilation Rate with zero shot.
 6. Discussion
+6.1. Error fixing by ZS4C
+To understand the errors that ZS4C is capable of fixing, we perform error analysis before and after running ZS4CGPT-3.5 on StatType-SO.
+To do so, we manually examine all the code snippets and their error logs returned by the compiler.
+We derived the following three types of errors for the code snippets of StatType-SO (Java code).
+Note that one code snippet could have multiple errors.
+  • Symbol Not Found: It occurs during compilation when the compiler encounters an identifier (such as a variable or method name) that it cannot recognize or resolve.
+  • Wrong Annotation: This error is triggered when an annotation is applied wrongly.
+  • Method Override/Implementation Error: This error happens when a method signature does not match the one in the superclass or interface, violating the rules of method overriding or implementation.
+We now show some examples of each type of error and how they were fixed by our approach.
+In Figure 5, the first code snippet (5(a)) demonstrates “Symbol Not Found” errors.
+Specifically, the ConfigParameters class extends ParameterHolder and uses ResourceSettings and Environment classes without defining them.
+This causes multiple “Symbol Not Found” errors because these classes and their respective methods and fields are unrecognized.
+In the second code snippet (5(b)), these errors are resolved by defining the missing classes ParameterHolder, ResourceSettings, and Environment within the same file.
+This ensures that all referenced classes are recognized and accessible during compilation, thus eliminating errors.
+public class xstream_class_34 {
+@XStreamAlias("config")
+public class ConfigParameters extends ParameterHolder {
+/* Error: ParameterHolder
+undefined */
+@XStreamImplicit(itemFieldName = "resource")
+private List<ResourceSettings> resources;
+/* Error: ResourceSettings
+undefined */
+@XStreamImplicit(itemFieldName = "env")
+private List<Environment> environments;
+/* Error: Environment
+undefined */
+// Other code
+(a) Code with errors (before running with ZS4C).
+public class xstream_class_34 {
+@XStreamAlias("config")
+public static class ConfigParameters extends ParameterHolder {
+/* Existing code (Figure a) */
+/* Missing classes are added below */
+public static class ParameterHolder { /* Fix: Added missing ParameterHolder class
+definition */
+public static class ResourceSettings { /* Fix: Added missing ResourceSettings class
+definition */
+private String name;
+public String getName() { return name; }
+public void setName(String name) { this.name = name; }
+public static class Environment { /* Fix: Added missing Environment class
+definition */
+private String name;
+public String getName() { return name; }
+public void setName(String name) { this.name is name; }
+(b) Code with fixes (after running with ZS4C).
+Figure 5. An example of “Symbol Not Found” error fixes using ZS4C for StatType-SO.
+In Figure 6, the first code snippet (6(a)) illustrates the “Wrong Annotation” errors.
+The annotation @Transactional is placed outside of any class, which causes the compiler to expect a class, interface, or enum declaration immediately following these annotations.
+This issue is addressed and fixed in the second code snippet (6(b)).
+The corrected placement of the annotations is directly before the GenericDaoImpl class inside the hibernate_class, ensuring proper association and adherence to Java syntax requirements.
+package hibernate;
+@Transactional
+@SuppressWarnings("unchecked")
+public class GenericDaoImpl<T, ID extends Serializable> implements GenericDao<T, ID> {
+// Other code
+(a) Code with errors (before running with ZS4C).
+package hibernate;
+public class hibernate_class { /* Fix: Correct annotation placement */
+@Transactional
+@SuppressWarnings("unchecked")
+public class GenericDaoImpl<T, ID extends Serializable> implements GenericDao<T, ID> {
+// Other code
+(b) Code with fixes (after running with ZS4C).
+Figure 6. An example of “Wrong Annotation” error fix using ZS4C for StatType-SO.
+In Figure 7, the first code snippet (7(a)) shows “Method Override/Implementation Error”.
+The class DetailDollarsConverter extends the superclass ReflectionConverter but fails to correctly override the marshal and unmarshal methods because the method signatures do not match those in the superclass, causing the @Override annotations to throw errors.
+These issues are resolved in the second code snippet (7(b)) by ZS4C.
+The corrected class DetailDollarsConverter now implements the Converter interface, ensuring the correct implementation of the marshal and unmarshal methods by modifying the code structure and method definitions.
+Table 9 presents the analysis performed on the StatType-SO dataset.
+The results indicate that ZS4C is highly effective in reducing various error types.
+This reduction is especially notable in the “Symbol Not Found” and “Method Override/Implement Error” categories, where every dataset showed a substantial decrease in errors.
+The results show the potential of ZS4C in other applications, such as fixing errors of code snippets that are generated by LLM and auto code completion in IDE (JetBrains Community, 2018; Pan et al., 2024).
+public class xstream_class_24 {
+public class DetailDollarsConverter extends ReflectionConverter {
+@Override /* Error: Method does not override
+marshal method in
+ReflectionConverter */
+public void marshal(Object obj, HierarchicalStreamWriter writer, MarshallingContext context) {
+super.marshal(obj, writer, context);
+writer.startNode("node4");
+writer.setValue(Double.toString(20));
+writer.endNode();
+@Override /* Error: Method does not override
+unmarshal method in
+ReflectionConverter */
+public Object unmarshal(HierarchicalStreamReader reader, UnmarshallingContext context) {
+return super.unmarshal(reader, context);
+// Other code
+(a) Code with errors (before running with ZS4C).
+public class xstream_class_24 {
+public class DetailDollarsConverter implements Converter { /* Correct interface implementation */
+private Mapper mapper;
+private ReflectionProvider reflectionProvider;
+public DetailDollarsConverter(Mapper mapper, ReflectionProvider reflectionProvider) {
+this.mapper = mapper;
+this.reflectionProvider = reflectionProvider;
+@Override /* Fixed by implementing Converter
+interface and defining marshal method
+as outlined below */
+public void marshal(Object obj, HierarchicalStreamWriter writer, MarshallingContext context) {
+ReflectionConverter reflectionConverter = new ReflectionConverter(mapper, reflectionProvider);
+reflectionConverter.marshal(obj, writer, context); /* Uses ReflectionConverter instance to
+marshal the object */
+writer.startNode("node4");
+writer.setValue(Double.toString(20));
+writer.endNode();
+/* Following the above similar
+logic, the ReflectionConverter instance
+is utilized to unmarshal the object. */
+// Other code
+(b) Code with fixes (after running with our ZS4C).
+Figure 7. Examples of “Method Override/Implementation” error fix using ZS4C for StatType-SO.
+Table 9. Error analysis before and after running ZS4CGPT-3.5 on StatType-SO.
+Before After
+Symbol
+Not Found
+Wrong
+Annotation
+Method
+Override/
+Implementation
+Error
+Symbol
+Not Found
+Wrong
+Annotation
+Method
+Override/
+Implementation
+Error
+Android 50 0 21 2 0 0
+JDK 21 0 3 0 0 0
+GWT 50 0 9 4 0 0
+Hibernate 50 1 2 17 0 1
+Joda-Time 50 0 0 0 0 0
+XStream 44 0 12 7 0 1
 6.2. Failure cases analysis
-To further understand the reasons that contribute to the failure of ZS4C in synthesizing compilable code even after applying Conversational Error Fixing, we conduct a qualitative analysis of the failed cases.
+To further understand the reasons that contribute to the failure of ZS4C with GPT-3.5 for StatType-SO dataset in synthesizing compilable code even after applying Conversational Error Fixing, we conduct a qualitative analysis of the failed cases.
 We collect all 33 failed cases and manually examine the conversation history between the compiler and LLM.
@@
 We encourage future research to investigate this direction.
+6.4. Cost analysis of API calls
+Table 10 summarizes the costs of ZS4C using the OpenAI API with GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4o) models along with StatType-SO and Python-SO datasets.
+The table compares the costs across one-shot, zero-shot, and three-shot prompt configurations for both models.
+As we can observe, it is economical to use both GPT-3.5 and GPT-4 to fix the code, which only costs 0.003 and 0.025 for one fix in the zero-shot setting, respectively.
+Using GPT-4 is more expensive than GPT-3.5, although it provides better performance.
+Meanwhile, the results of RQ4 indicate that zero-shot achieves comparable performance as one-shot and three-shot prompt settings.
+This finding is practically significant, as zero-shot prompt saves the cost of retrieving examples and reducing the prompt size, while still achieving comparable performance.
+Table 10. Cost analysis of OpenAI API for various prompt settings based on input token usage.
+StatType-SO Python-SO
+Total Cost (USD) Cost/case (USD) Total Cost (USD) Cost/case (USD)
+One-shot (GPT-3.5) \$1.143 \$0.004 \$1.635 \$0.003
+Zero-shot (GPT-3.5) \$0.706 \$0.003 \$0.954, \$0.002
+Three-shot (GPT-4) \$19.556 \$0.073 \$28.609 \$0.053
+One-shot (GPT-4) \$10.92 \$0.041 \$15.826 \$0.029
+Zero-shot (GPT-4) \$6.596 \$0.025 \$9.358 \$0.017
+Total \$38.921 \$0.146 \$56.382 \$0.104
 6.5. Threats to validity
 Threats to external validity.
-Consistent with prior studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), we utilize the StatType-SO benchmark as the evaluation dataset.
-This dataset, comprising manually verified code snippets from six prominent Java libraries across diverse domains, may not be representative of libraries in all programming languages.
-Nonetheless, ZS4C is language-agnostic and it does not rely on Java-specific constructs.
-We have chosen ChatGPT (GPT-3.5-Turbo) as the base LLM for our study.
-Consequently, our findings might differ if alternative LLMs, such as GPT-4, are employed.
+Consistent with prior studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019), we utilize the StatType-SO benchmark and a newly constructed Python dataset as the evaluation datasets.
+Those two datasets comprising manually verified code snippets across diverse domains, may not be representative of libraries in all programming languages.
+Nonetheless, ZS4C is language-agnostic and can be applied to any programming language.
+Another external threat is that the code snippets in our studied dataset are often small (with an average of 14-82 LOC across libraries), so our findings might not be generalized to larger code snippets.
+In this study, we have chosen GPT-3.5 and GPT-4 as our base LLMs for our study.
+Consequently, our findings might differ if alternative LLMs, such as Gemini, are employed.
+We encourage future research on more datasets with larger code snippets and more base LLMs.
 Threats to internal validity.
@@
 Given the context window token constraints of ChatGPT, we consider adding up to one demonstration example to ZS4C and form its one-shot version, as incorporating more examples might breach the token limit.
-Furthermore, only the most recent chat history is retained in ConversationFixing once the context window limit is approached.
+Furthermore, only the most recent chat history is retained in ConversationFixing once the context window limit is approached for GPT-3.5.
+As introduced in Section 3.2, the same APIs could appear in different packages (i.e., API collision) and make it challenging to decide which packages to use during Conversational Error Fixing.
+This could be an internal threat to the validity.
+To mitigate the threat, we selected the one that satisfies pre-defined constraints (i.e., invoked functions and attributes) following the previous study (Subramanian et al., 2014).
+If there exist multiple libraries available after resolving constraints, we select the one that was released most recently.
+Through such process, we achieved a low API collision rate 5.5% (82 out of 1,494 cases exhibit API collisions) on StatType-SO dataset.
 7. Conclusion